### v2

# Two dimensional dimers beyond planarity

Abstract
We consider a generalisation of the double dimer model which includes several models of interest, such as the monomer double dimer model, the dimer model, the Spin O(N) model, and it is related to the loop O(N) model. We prove that on two-dimension like graphs (such as slabs), both the correlation function and the probability that a loop visits two vertices converge to zero as the distance between such vertices gets large. Our analysis is by introducing a new (complex) spin representation for all models belonging to this class, and by deriving a new proof of the Mermin-Wagner theorem which does not require the positivity of the Gibbs measure. Even for the well studied dimer and double dimer model our results are new since - not relying on solvability and Kasteleyn’s theorem - they hold beyond the framework of planar graphs.
1 Introduction
The dimer model is the study of the set of perfect matchings of a graph. It attracts interest from a wide range of perspectives, including probability, combinatorics, statistical mechanics, graph theory. The double dimer model consists in the superposition of two perfect matchings of a graph drawn independently at random with uniform distribution, its realisations can be viewed as collections of disjoint self-avoiding loops.
The rigorous study of such models achieved a breakthrough with the seminal works of Kasteleyn [15], Temperley and Fisher [31], who proved a theorem (known as Kasteleyn’s theorem) allowing an exact computation of the number of dimer configurations on planar graphs. Understanding the properties of perfect matchings beyond the framework of Kasteleyn’s theorem is much harder and of great interest. New results in this direction have recently been achieved in the study of the dimer problem in $\mathbb{Z}^{d}$ for $d\geq 3$ [6, 28, 30, 27] and on weakly planar graphs [19].
The present paper considers the dimer model, the double dimer model and a generalisation of such models on two dimensional (not necessarily planar) graphs. We will discuss the most general setting in Theorem 7.2 and Section 7.1 below, will now focus our discussion on the monomer double dimer model, a special case of our general framework which is of central interest. The configurations of the monomer double dimer model are obtained by selecting a random set of vertices (monomers) and then drawing at random two independent sub-graphs such that the degree of each monomer is zero and the degree of the remaining vertices is precisely one. The realisations of such a model can be viewed as disjoint collections of self-avoiding loops (see the left of Figure 1.1 for a configuration with four monomers as an example, where the edges of the two sub-graphs are coloured by blue and red respectively). The model depends on a parameter, the monomer activity, which rewards the total number of monomers. When the monomer activity is zero no monomer is present and the model reduces to the double dimer model.
Figure 1.1: Left: A configuration of the monomer double dimer model in a box of $\mathbb{Z}^{2}$. Right: A dimer configuration on the duplicated graph. The red, orange and blue dimers represent the vertical dimers and the dimers of the lower and upper box, respectively. Projecting the upper and lower box onto the same box leads to the configuration on the left.
If the monomer activity is positive, it is not necessarily the case that each vertex is touched by a loop. Alternatively, the configurations of this model on a graph $G$ can be viewed as a perfect matching of a graph corresponding to two copies of $G$, with each vertex of the first copy being connected by an edge to the corresponding vertex of the second copy. The monomers in the first formulation then correspond to dimers on the edges connecting the two copies in the second formulation, as it is represented in Figure 1.1.
The presence a positive density of monomers makes the model computationally intractable [11] and Kasteleyn’s theorem inapplicable, understanding simple properties of this basic model is then of great interest.
We state our results on two dimensional slabs with arbitrary width (the width may even ‘slowly’ grow to infinity with the number of vertices), the two-dimensional planar torus is then a special case of our theorems.
Our first main theorem (Theorem 1.1 below) provides an upper bound for the typical length of the loops in the monomer double dimer model. Our theorem states that, for each choice of the monomer activity, the expected loop length divided by the number of vertices in the graph converges to zero as the number of vertices in the graph tend to infinity. Our result then implies that in two dimensional graphs the probability that two vertices are connected by the same loop decays to zero with their distance, contrary to dimensions three and higher [27].
Our second main theorem (Theorem 1.2 below) involves monomer-monomer correlations in the dimer model, namely the ratio between the number of perfect matchings of a graph from which two vertices (and all the edges connected to them) have been removed and the number of perfect matchings of the graph. We show that the Cesaro sum of monomer-momomer correlations converges to zero as the number of vertices in the graph tend to infinity.
Notice that our first main theorem is known in the special case of zero monomer activity (the double dimer model) on planar graphs [9, 22] and our second main theorem is known for the dimer model on planar graphs [8]. The novelty of our work is to prove such results beyond the framework of Kasteleyn’s theorem, e.g. on two dimensional non-planar graphs and/or in the presence of a positive density of monomers. Indeed, even on planar graphs, the dimer model with a positive density of monomers is generally not solvable. Only if the monomers are present on the boundary of a bounded domain Kasteleyn’s theorem is applicable, since in this case one can map the system into a dimer model in an enlarged graph (see [2, 18] for recent works).
Our results hold in even more general frameworks. For example, they hold for the multiple dimer model, with dimers having $2N$ possible ‘colours’, such that the number of dimers of even colour $i$ touching any vertex is arbitrary and equals the number dimers of colour $i+1$ touching the same vertex. The monomer double dimer model then corresponds to the special case $N=1$, with the number of dimers of colour $1$ touching each vertex being equal to the number of dimers of colour $2$ touching the same vertex and this number being either 00 or $1$. We treat such generalisations in Theorem 7.2 and in Section 7.1 below. Such a general framework also includes a model which is closely related to the loop O(N) model (when $N$ is even), which has been mostly studied on the hexagonal lattice (see e.g. [10, 26] and references therein). The (minor) difference between our model and the loop O(N) model on the hexagonal lattice is that in the latter one no loop of length two is present. Our techniques are complementary to those in [10], since they provide a lower bound for the probability of observing a long loop, and, contrary to [10], our method can be applied only for integer values of $N$, but for any value of the inverse temperature parameter and even on graphs which are not necessarily hexagonal (and which are symmetric enough to employ the reflection positivity method, for example periodic boundary conditions are a necessary ingredient).
A key ingredient of our proofs is a spin representation relating connection probabilities and monomer correlations in the monomer double dimer model to spin-spin correlations. Our spin representation presents a complex duality relation between the two spin components which plays a crucial role in our analysis and makes the representation invariant with respect to global spin rotations. The main obstacle in using such representation is that, since the measure is non positive (and complex), one cannot employ the classical strategies [7, 14, 17, 25] based on the spin-wave deformation, or directly apply the older approach of [13]. Indeed, these approaches use the Cauchy-Schwarz inequality in an essential way, which does not produce useful bounds for non-positive measures. We overcome these obstacles introducing a new proof of a Mermin-Wagner theorem for such complex measures. We rely on the reflection positivity property of such models, which allows us to derive a weak form of a Cauchy-Schwarz inequality for a restricted (but useful) class of functions (see Theorem 6.3 below). This inequality is used to show that the magnetisation vanishes if we take the limit of small external fields appropriately. A perturbative analysis allows us to show that a vanishing magnetisation implies that the Cesaro sum of two point functions is small (see Proposition 5.4 below). Our general strategy for proving the Mermin Wagner theorem is new and can then be applied to any spin system with a non necessarily positive (but reflection positive) measure.
Parallel to our work, a related complex spin representation has been introduced in [3] for the derivation of infrared bounds. The results which are presented in [3] have been obtained independently from us.
Main results.
Consider a finite undirected bipartite graph $G=(V,E)$. A dimer configuration is a spanning sub-graph of $G$ such that every vertex has degree one. We let $\mathcal{D}_{G}$ be the set of dimer configurations in $G$. Given a set $A\subset V$, we let $G_{A}$ be the subgraph of $G$ with vertex set $V\setminus A$ and with edge set consisting of all the edges in $E$ which do not touch any vertex in $A$. We let $\mathcal{D}_{G}(A)$ be the set of dimer configurations in $G_{A}$.
The configuration space of the monomer double dimer model is denoted by $\Omega$ and it corresponds to the set of triplets $\omega=(M,d_{1},d_{2})$ such that $M\subset V$ and $(d_{1},d_{2})\in\mathcal{D}_{G}(M)\times\mathcal{D}_{G}(M)$. We refer to the first element of the triplet $\omega$ as a set of monomers, and to the second and third element as set of dimers. For example, a configuration with four monomers in the box of $\mathbb{Z}^{2}$ can be viewed in the left of Figure 1.1. As one can see from the figure, any configuration $(M,d_{1},d_{2})\in\Omega$ can be viewed as a collection of disjoint self-avoiding loops, with loops possibly having length two and corresponding to the superposition of two dimers on the same edge. We let $\mathcal{M}:\Omega\mapsto V$ be the random variable corresponding to the set of monomers, i.e, for each $\omega=(M,d_{1},d_{2})\in\Omega$ $\mathcal{M}(\omega):=M$. We define a probability measure on $\Omega$,
$\forall\omega\in\Omega\quad\mathbb{P}_{G,N,\rho}(\omega):=\frac{\rho^{|\mathcal{M}(\omega)|}\,N^{\mathcal{L}(\omega)}}{\mathbb{Z}_{G,N,\rho}},$ (1.1)
where $\rho,N\geq 0$ are some parameters, $\mathcal{L}(\omega)$ is the number of loops in $\omega$ and $\mathbb{Z}_{G,N,\rho}$ is a normalisation constant. We refer to $\rho$ as the monomer activity and use $\mathbb{E}_{G,N,\rho}$ to denote the expectation with respect to (1.1). The monomer double dimer model illustrated in Figure 1.1 corresponds to the special case of $N=1$.
We define $L_{x}={L}_{x}(\omega)$ as the loop which touches the vertex $x\in V$ in the configuration $\omega\in\Omega$ and set $L_{x}(\omega)=\emptyset$ if the configuration $\omega$ has a monomer at $x$ Moreover, we denote by $|{L}_{x}|=|{L}_{x}|(\omega)$ the length of the loop $L_{x}(\omega)$, namely the number of vertices which are touched by such a loop, and set $|{L}_{x}|(\omega)=0$ if $\omega$ has a monomer at $x$.
We let $\mathbb{S}_{L,K}$ denote the two dimensional slab torus, which we identify with the set $\big{(}(-\frac{L}{2},\frac{L}{2}]\cap\mathbb{Z}\big{)}\times\big{(}(-\frac{L}{2},\frac{L}{2}]\cap\mathbb{Z}\big{)}\times\big{(}(-\frac{K}{2},\frac{K}{2}]\cap\mathbb{Z}\big{)}$. In the special case $K=1$, $\mathbb{S}_{L,1}$ corresponds to the planar torus, while $\mathbb{S}_{1,1}$ is a single point. We use $o\in\mathbb{S}_{L,K}$ to denote the origin. In (1.1) and in the corresponding expectation we use the subscript ${}_{L,K}$ rather than ${}_{\mathbb{S}_{L,K}}$ in order to simplify the notation. Our theorem states that the expected loop length divided by the volume converges to zero in the limit of large slabs.
Theorem 1.1.
Suppose that $\rho>0$ and $N\in\mathbb{N}$ or that $\rho=0$ and $N=1$. There exists $c=c(\rho,N)<\infty$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$ satisfying $K\leq\sqrt{\log(L)}$,
$\frac{\mathbb{E}_{L,K,N,\rho}\big{(}|L_{o}|\big{)}}{|\mathbb{S}_{L,K}|}\leq c\sqrt{\frac{K}{\log L}}.$ (1.2)
Since $\mathbb{E}_{L,K,N,\rho}\big{(}|L_{o}|\big{)}=\sum_{x\in\mathbb{S}_{L,K}}\mathbb{P}_{L,K,N,\rho}(o\leftrightarrow x)$, our theorem implies that the probability that two distant vertices are visited by the same loop is not uniformly positive, contrary to dimension three and higher and small monomer activity [28].
Our previous theorem can also be stated for other quantities of interest in the monomer double dimer model with arbitrary monomer activity, such as the Cesaro sum of the two-point function $\mathcal{G}^{1}_{G}(x,y)$ which is introduced in equation (3.5) below. In the special case of zero monomer activity such a two-point function reduces to the monomer-monomer correlation in the dimer model,
$\forall x,y\in V\quad\mathcal{C}_{G}(x,y):=\frac{|\mathcal{D}_{G}(\{x,y\})|}{|\mathcal{D}_{G}|},$ (1.3)
a quantity of central interest in the study of this model (if $G=\mathbb{S}_{L,K}$, we use the subscript ${}_{L,K}$ rather than ${}_{\mathbb{S}_{L,K}}$, in order to simplify the notation). Our third main theorem provides a upper bound on the Cesaro sum of such a function on two dimensional slabs.
Theorem 1.2.
There exists $c<\infty$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$ satisfying $K\leq\sqrt{\log(L)}$, we have
$\sum\limits_{x\in\mathbb{S}_{L,K}}\frac{\mathcal{C}_{L,K}(o,x)}{|\mathbb{S}_{L,K}|}\leq c\sqrt{\frac{K}{\log L}}$ (1.4)
Contrary to (1.4) the Cesaro sum of monomer-monomer correlations is known to be uniformly positive in the $d$-dimensional torus if $d\geq 3$ [30]. Note that both our theorems also hold for the planar torus, corresponding to the case $K=1$, and even in this case Theorem 1.1 is new when $\rho>0$, being Kasteleyn’s theorem inapplicable.
Discussion.
When $N=1$ the monomer double dimer model (and the general class of models introduced in Section 3 below) is expected to undergo a Kosterlitz-Thouless phase transition with respect to the variation of the monomer activity on two dimensional graphs. This means that the connection probability exhibits exponential decay for large enough monomer activity (which has been proved in [4]) and polynomial decay for small enough (but possibly positive) monomer activity. We may hope that the spin representation may help us to import techniques from spin systems in the framework of perfect matchings to make further progress. For example, one may hope to apply the recent methods of [1, 12] or the technique of [16] to show that the connection probability decays not faster than polynomially on two dimensional graphs if the monomer activity is small enough and $N=1$, or to make further progress with the dimer problem (with or without monomers) in dimensions three and higher [28, 30].
Paper organization.
In Section 2 we introduce the complex spin representation and state our main theorem for spins. In Section 3 we introduce our generalized double dimer model, discuss some special cases and make a connection between dimers and spins. In Section 4 we introduce the reflection positivity property and discuss its consequences. In Section 5 we present some probabilistic estimates, introduce a central quantity, the magnetisation, and present a perturbative analysis in the limit of vanishing external field which allows us to lower bound the magnetisation by the Cesaro sum of two point functions at zero external field. Section 6 is the central part of our paper, here we introduce our central tool, a weak version of the Cauchy Schwarz for our complex measure. In Section 7 we present the proof of our main theorems and describe the generalisation to arbitrary number of colours. Finally, in Section 8 we present the proof of our proposition which relates spins and dimers.
Notation.
Unless otherwise specified, the constants $c$, $c_{i}$, $i=1,2,\ldots$, appearing in the text will always be positive, finite and will possibly depend on the parameter $\beta$, which is introduced below and corresponds to the inverse monomer activity in the monomer double dimer model. Their value may change from line to line. Any further dependence of such constants from the other parameters will be specified explicitly. In the whole paper the graph $G=(V,E)$ will always be a undirected finite bipartite graph. Moreover, $d(x,y)$ will be used to denote the graph distance. We use the notation $\mathbb{N}_{0}=\{0,1,\ldots\}$, =$\mathbb{N}=\{1,2,\ldots\}$, $\mathbb{R}^{+}_{0}=\{x\in\mathbb{R}:x\geq 0\}$. Finally, we use the abbreviations ‘LHS’ for ‘left-hand side’ and ‘RHS’ for ‘right-hand side’.
2 The spin representation
We let $G=(V,E)$ be a undirected finite bipartite graph with $o\in V$ a prescribed vertex. We let $V^{e}$ (resp. $V^{o}$) be the set of vertices having even (resp. odd) graph distance from $o$. We let $U:\mathbb{N}_{0}\mapsto\mathbb{R}^{+}_{0}$ be a weight function, which defines our model. In the whole paper we assume that the non negative sequence $\big{(}U(n)\big{)}_{n\in\mathbb{N}_{0}}$ is sumable, that $U(n)>0$ for some $n>0$, and that $U(n)\leq 1$ for any $n\in\mathbb{N}_{0}$. We introduce a ‘local’ spin space $S=[0,2\pi)^{2}$. We define for each $s=(s^{1},s^{2})\in S$ the function
$g(s):=\frac{1}{(2\pi)^{2}}\sum\limits_{n=0}^{\infty}U(n)e^{-n\,i\,(s^{1}-s^{2})}.$ (2.1)
We denote by $\Omega_{s}:=S^{V}$ the configuration space and define for each $\boldsymbol{s}=(s_{z})_{z\in V}\in\Omega_{s}$, with $s_{z}=(s_{z}^{1},s_{z}^{2})$,
$\boldsymbol{g}(\boldsymbol{s}):=\prod_{i\in V^{e}}g(s_{i})\prod_{j\in V^{o}}\overline{g}(s_{j})$ (2.2)
and interpret $s^{1}$, $s^{2}$ as the angles of the two components of the spin. We define the function
$H(\boldsymbol{s})$ :=$:=\sum\limits_{\{i,j\}\in E}e^{i(s_{i}^{1}-s_{j}^{1})}+\sum\limits_{\{i,j\}\in E}e^{-i(s_{i}^{2}-s_{j}^{2})}+\,h\sum\limits_{k\in V}\,2\,\cos(rs^{1}_{k}).$ (2.3)
In the previous expressions (and in the whole paper) we use the convention that the sum over indirected edges $\{i,j\}$ the first vertex, $i$, is always even and the second vertex, $j$, is always odd. Moreover, $h,r\geq 0$ are some parameters. We define the complex measures assigning to each measurable complex function $f:\Omega_{s}\mapsto\mathbb{C}$ the weight
$\langle f\rangle_{G,\beta,h,r}$ $:=\frac{1}{Z_{G,\beta,h,r}}\int_{\Omega_{s}}\boldsymbol{ds}\,\boldsymbol{g}(\boldsymbol{s})\,\,e^{\beta H(\boldsymbol{s})}\,\,f(\boldsymbol{s})$ (2.4)
$Z_{G,\beta,h,r}$ $:=\int_{\Omega_{s}}\boldsymbol{ds}\,\boldsymbol{g}(\boldsymbol{s})\,\,e^{\beta H(\boldsymbol{s})},$ (2.5)
where we integrate with respect to a product of Lebeasgue measures in $S^{V}$ and $\beta\geq 0$ is a parameter, the inverse temperature, which will play the role of the inverse monomer activity in the monomer double dimer model. Clearly, if $h=0$ the operator $\langle\,\,\rangle_{G,\beta,h,r}$ does not depend on the value of $r$.
Our spin representation is invariant with respect to global spin rotations. Indeed, due to our choice of (2.1), if we add the same constant to the ‘angles’ $s_{x}^{1}$ and $s_{x}^{2}$ for any $x\in V$ the integrand in (2.4) does not change. This feature will turn out to be essential for the proof of our main theorem, and seems to be the analogous of the invariance under rotation property in Spin O(N) models with $N\geq 2$.
Our ultimate goal is to obtain information on the behaviour of the two-point function when $h=0$. We consider two different two-point functions, corresponding to the cases $\ell=1$ and $\ell=2$ in equation (2.6) below. These are related to two different important quantities in our general double dimer model (we explore this connection in Proposition 3.3 below). We will obtain information on such two point functions by considering the limit $h\rightarrow 0$ of our spin system when $r=1$ and $r=2$ in (2.3) respectively. We now state our main theorem about decay of correlations in the spin system (2.4). When the graph $G$ is the slab torus $\mathbb{S}_{L,K}$, we use the subscript ${}_{L,K}$ rather than ${}_{\mathbb{S}_{L,K}}$ in order to simplify the notation.
Theorem 2.1.
Fix $\ell\in\{1,2\}$, let $\beta\geq 0$ be arbitrary. There exists $c=c(\beta)\in(0,\infty)$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$ satisfying $K\leq\sqrt{\log(L)}$, we have
$0\leq\sum\limits_{x\in\mathbb{S}_{L,K}}\frac{1}{|\mathbb{S}_{L,K}|}\langle\,\cos(\ell\,s_{o}^{1})\cos(\ell\,s_{x}^{1})\,\rangle_{L,K,\beta,0,0}\leq c\sqrt{\frac{K}{\log L}}$ (2.6)
Most of the paper is devoted to the proof of this theorem, from which we deduce Theorems 1.1 and 1.2. Its proof is presented in Section 7 below.
3 A generalized double dimer model
We introduce a generalisation of the double dimer model, in which each vertex is touched by the same number of dimers for each type ($1$=blue and $2=$red) and this number is an arbitrary non-negative integer. This generalisation includes the monomer double dimer model, in which such a number is either zero or one, and the double dimer model, in which such a number is precisely one. In Section 7.1 below we deal with the case of arbitrary number of colours.
3.1 Definition
Graph. We assume that $r\in\mathbb{N}$. In order to define a generalized double dimer model corresponding to the spin system on the bipartite graph $G=(V,E)$, we need to enlarge the graph, by introducing a ghost vertex $g$, representing the effect of the external magnetic field, and a source vector, $s$, which allows us to represent correlation functions for the spin system in terms of dimers. Each such vertex is connected by an edge to a vertex of the original graph $G$. The enlarged graph is denoted by $G_{en}:=(V_{en},E_{en})$, with $V_{en}:=V\cup\{g,s\}$, $s\neq g$, $E_{en}:=E\cup E_{g}\cup E_{s}$, where $E_{g}:=\big{\{}\{x,g\}\,:\,x\in V\big{\}}$ and $E_{s}:=\big{\{}\{x,s\}\,:\,x\in V\big{\}}$. We write $x\stackrel{{\scriptstyle en}}{{\sim}}y$ for $\{x,y\}\in E_{en}$ and $x\sim y$ for $\{x,y\}\in E$. We recall that $V^{e}\subset V$ (resp. $V^{o}\subset V$) is the set of vertices having even (resp. odd) graph distance from the origin $o\in V$ in $G\subset G_{en}$.
Configurations. System realisations are triplets $w=(m^{1},m^{2},\pi)$, which are now defined. We first define the pairs $m=(m^{1},m^{2})\in\Sigma$, where
$\Sigma:=\big{\{}(m^{1},m^{2})\in\mathbb{N}_{0}^{E_{en}}\times\mathbb{N}_{0}^{E_{en}}\,:\,\sum_{y\stackrel{{\scriptstyle en}}{{\sim}}x}m^{1}_{\{x,y\}}=\sum_{y\stackrel{{\scriptstyle en}}{{\sim}}x}m^{2}_{\{x,y\}}\,\,\mbox{ and }\,\,m_{\{x,g\}}^{i}\in r\mathbb{N}_{0}\,\,\,\forall x\in V,\forall i\in\{1,2\}\big{\}}.$
We interpret for each edge $e\in E_{en}$, $m^{i}_{e}$ as the number of dimers of colour $i\in\{1,2\}$ ($i$-dimers) on $e$. By definition, each vertex in $V$ can only be touched by the same number of blue and red dimers, such a constraint however does not hold for $s$ or $g$. Moreover, the number of dimers on each edge touching the ghost vertex is only allowed to be a multiple of $r$.
Given $m\in\Sigma$, we let $\mathcal{P}(m)$ be the set of pairing functions $\pi=(\pi_{x})_{x\in V}$ such that, at each odd vertex $x\in V^{o}$, $\pi_{x}$ is a bijection from the set of blue dimers touching $x$ to the set of red dimers touching $x$, and at each even vertex $x\in V^{e}$, $\pi_{x}$ is a bijection from the set of red dimers touching $x$ to the set of blue dimers touching $x$. Note that pairing functions are not associated to $g$ or $s$, hence the dimers touching these vertices are always ‘left unpaired’.
Hence, any triplet $w=(m^{1},m^{2},\pi)$ with $m=(m^{1},m^{2})\in\Sigma$ and $\pi\in\mathcal{P}(m)$ can be viewed as a collection of paths, where each path is a sequence of dimers with alternating colours. Each dimer in such a sequence is the image and it is mapped to a dimer of a different colour through the pairing function. Such a path can be ‘closed’ and lying completely in $G$ or ‘open’ and then start and end at a ghost or source vertex. The ghost and the source vertex can then be viewed as sources of open paths. We will refer to the closed paths as loop and to the open ones as walk.
We let $\mathcal{W}$ be the set of configurations $w=(m,\pi)=(m^{1},m^{2},\pi)$ such that $m=(m^{1},m^{2})\in\Sigma$ and $\pi\in\mathcal{P}(m)$.
Local time and source vector. Given $m=(m^{1},m^{2})\in\Sigma$, we define the local time at each vertex $x\in V$ as
$n_{x}(m)=\sum_{\begin{subarray}{c}y\stackrel{{\scriptstyle e}}{{\sim}}x\end{subarray}}m^{1}_{\{x,y\}}=\sum_{\begin{subarray}{c}y\stackrel{{\scriptstyle e}}{{\sim}}x\end{subarray}}m^{2}_{\{x,y\}},$ (3.1)
corresponding to the total number of blue or red dimers touching that vertex. Given an element $m\in\Sigma$, we let $\partial m=\big{(}(\partial m)^{1}_{x},(\partial m)^{2}_{x}\big{)}_{x\in V}\in\mathbb{Z}^{V\times V}$ be the source vector, where
$(\partial m)^{i}_{x}:=m^{i}_{\{x,s\}}\in\mathbb{N}_{0}$
corresponds to the number of $i$-dimers on $\{x,s\}$. Note that these quantities are not defined for $x=g$ or $x=s$. By a slight abuse of notation sometimes we will use $n_{x}(w)$ and $(\partial w)^{i}_{x}$ rather than $n_{x}(m)$ and $(\partial m)^{i}_{x}$ for $w=(m,\pi)\in\mathcal{W}$, and we consider $n_{x}$ and $(\partial w)^{i}_{x}$ them as random variables in $\mathcal{W}$. We also use $m^{i}_{e}(w)$ for the random variable corresponding to the number of $i$ dimers on $e$.
Measures and two-point functions. We define the (unnormalised, positive) measure in $\mathcal{W}$ assigning to each $w=(m,\pi)\in\mathcal{W}$ the weight
$\mu_{\beta,h,r}(w):=\tilde{\mu}_{\beta,h,r}(m)\prod_{x\in V}\frac{U\big{(}k_{x}(w)\big{)}}{n_{x}(w)!},$ (3.2)
where $k_{x}(w):=n_{x}(w)-m_{\{x,g\}}^{2}(w)$ for each $x\in V$, and
$\tilde{\mu}_{\beta,h,r}(m):=\prod_{e\in E}\frac{\beta^{m_{e}^{1}}}{m^{1}_{e}!}\frac{\beta^{m_{e}^{2}}}{m^{2}_{e}!}\prod_{x\in V}\frac{(\beta h)^{\frac{m_{\{x,g\}}^{1}}{r}}}{\left(\frac{m^{1}_{\{x,g\}}}{r}\right)!}\frac{(\beta h)^{\frac{m_{\{x,g\}}^{2}}{r}}}{\left(\frac{m^{2}_{\{x,g\}}}{r}\right)!}.$
Note that the weight assigned by $\mu_{\beta,h,r}(w)$ does not depend on $\pi$, but only on the number of dimers on each edge and on their colour. We use $\mu_{\beta,h,r}$ also to denote expectations, hence, given $f:\mathcal{W}\mapsto\mathbb{R}$, we define
$\mu_{\beta,h,r}(f):=\sum\limits_{w\in\mathcal{W}}\mu_{\beta,h,r}(w)f(w).$
We now define the partition function of the random walk loop soup with no sources
$\mathcal{Z}_{G,\beta,h,r}:=\mu_{\beta,h,r}(\{\partial m=(0,0)\}),$ (3.3)
a probability measure on the space of paths with no sources $\mathcal{W}^{\ell}:=\{w\in\mathcal{W}\,:\,\partial w=(0,0)\}$,
$\forall w\in\mathcal{W}^{\ell}\quad\mathcal{P}_{G,\beta,h,r}(w):=\frac{\mu_{\beta,h,r}(w)}{\mathcal{Z}_{G,\beta,h,r}},$ (3.4)
the expectation with respect to (3.4), $\mathcal{E}_{G,\beta,h,r}$, and the two-point functions
$\mathcal{G}_{G,\beta}^{(1)}(x,y):=\frac{\mu_{G,\beta,0,0}(x\rightarrow y)}{\mathcal{Z}_{G,\beta,0,0}}\quad\quad\quad\mathcal{G}_{G,\beta}^{(2)}(x,y):=\mathcal{P}_{G,\beta,0,0}(x\leftrightarrow y),$ (3.5)
where $\{x\leftrightarrow y\}\subset\mathcal{W}^{\ell}$ is the event that there exists a loop visiting both $x$ and $y$ and $\{x\rightarrow y\}\subset\mathcal{W}$ is the event that there exists an walk ending with a dimer on $\{x,s\}$ and a dimer on $\{y,s\}$ and all the remaining paths are loops. Note that these quantities are defined under the assumption $h=0$, hence possible walks are not allowed to end at the ghost vertex.
3.2 Special cases
Some choices of the weight function lead to well known models in statistical mechanics, combinatorics, probability theory, and physics. We will explain these connections in the most interesting case $h=0$, in which case the measure (3.2) takes a particularly simple form, with no dimer touching the ghost vertex almost surely and open paths only terminating at the source vertex when present. We will keep assuming that $G=(V,E)$ is bipartite.
Monomer double dimer model.
We obtain the monomer double dimer model when $U(0)=U(1)=1$ and $U(n)=0$ for all $n>1$. Under this choice only configurations such that each vertex is either touched by precisely one blue and one red dimer or by no dimer are allowed. We let $\tilde{\Sigma}:=\big{\{}(m^{1},m^{2})\in\{0,1\}^{E}\times\{0,1\}^{E}:\sum_{y\sim x}m_{\{x,y\}}^{1}=\sum_{y\sim x}m_{\{x,y\}}^{2}\in\{0,1\},\forall x\in V\big{\}}.$ The partition function (3.3) then takes the simple form
$\mathcal{Z}_{G,\beta,0,0}=\sum\limits_{(m^{1},m^{2})\in\tilde{\Sigma}}\prod_{e\in E}\beta^{m^{1}_{e}+m^{2}_{e}}=\beta^{|V|}\,\mathbb{Z}_{G,1,\frac{1}{\beta}},$
where $\mathbb{Z}_{G,1,\frac{1}{\beta}}$ was defined in (1.1). For the first identity above we used that $h=0$ and the fact that, given a realisation $m\in\Sigma$ such that each vertex $x\in V$ is either touched by precisely one red and blue dimer or is left untouched, $\mathcal{P}(m)$ contains just one element. For the second identity we observed that the number of monomers plus the total number of dimers always equal $|V|$. Hence we deduce that the probability measure (3.4) then corresponds to the probability measure in $\tilde{\Sigma}$ assigning to each $m\in\tilde{\Sigma}$ the weight
$\frac{\prod_{e\in E}\beta^{m_{e}^{1}+m_{e}^{2}}}{\mathcal{Z}_{G,\beta,0,0}}.$
and that, for each pair of vertices $x,y\in V$,
$\mathcal{G}^{(2)}_{G,\beta}(x,y)=\mathbb{P}_{G,1,\frac{1}{\beta}}(x\leftrightarrow y),$ (3.6)
where we recall the definition (1.1) and, by a slight abuse of notation, we use $\{x\leftrightarrow y\}\subset\Omega$ also to denote the event that there exists a loop visiting both $x$ and $y$ in the monomer double dimer model.
Recalling the description of the monomer double dimer model as a dimer model on the duplicated torus, the two-point function $\mathcal{G}^{(1)}_{G,\beta}(x,y)$ can be viewed as the ratio between the weight of all dimer configurations on the duplicated torus with monomers displaced at two vertices of such a graph and the weight of all configurations with no such monomers.
Double dimer model and perfect matchings.
We obtain the double dimer model (corresponding to the monomer double dimer model with zero monomer activity) if $U(1)=1$ and $U(n)=0$ for each $n\neq 0$, in other words each vertex is touched by precisely one blue and red dimer. The function $g(s)$ defining the corresponding spin measure (2.4) is then
$g(s)=\frac{1}{(2\pi)^{2}}e^{-i(s^{1}-s^{2})}.$
Taking $\beta=1$ we have that the partition function (3.3) takes the simple form
$\mathcal{Z}_{G,1,0,0}=|\,\mathcal{D}_{G}(\emptyset)|^{2}.$
We further have that
$\mu_{G,1,0,0}(x\rightarrow y)=|\,\mathcal{D}_{G}(\emptyset)|\,|\,\mathcal{D}_{G}(\{x,y\})|$
(see also [30, Figure 2]). Hence from the definition (3.5) and the two previous expressions we obtain the following identity, which is key for the proof of Theorem 1.2
$\mathcal{G}_{G,1}^{(1)}(x,y)=\frac{|\,\mathcal{D}(\{x,y\})|}{|\,\mathcal{D}(\emptyset)|}$ (3.7)
and holds for any pair of distinct vertices $x,y\in V$.
The XY model.
We obtain the double dimer representation of the XY model when $U(n)=1$ for each $n\in\mathbb{N}_{0}$. In this case the random loop model (2.4) is equivalent to the ones which have been defined in [12, 23] (to see this connection, one should interpret the two colours of the dimers as the two possible directions of the edges in the multigraph which is considered in [12]), where it is proved that such models are a representation of the XY model.
3.3 From spins to double dimers
The next proposition states a correspondence between observables in our generalised double dimer model and in the spin system which was defined in the Section 2. For each $z\in V$, $u\in\mathbb{N}_{0}$, we define the functions $g_{u,z}:S\mapsto\mathbb{C}$, $g_{z}:S\mapsto\mathbb{C}$, which are
$g_{z}(s)=\begin{cases}\frac{U(0)e^{-2h\cos(rs_{o}^{1})}}{g(s)}&\mbox{ if $z\in V^{e}$}\\ \frac{U(0)e^{-2h\cos(rs_{o}^{1})}}{\overline{g}(s)}&\mbox{ if $z\in V^{o}$}\end{cases},\quad\quad g_{u,z}(s)=\begin{cases}\frac{U(u)}{g(s)}e^{-ui(s^{1}-s^{2})}&\mbox{ if $z\in V^{e}$}\\ \frac{U(u)}{\overline{g}(s)}e^{ui(s^{1}-s^{2})}&\mbox{ if $z\in V^{o}$}\end{cases}.$ (3.8)
The idea is that taking the average of the function $e^{n\,i\,s_{x}^{i}}$, $x\in V$, imposes the presence of $n$ dimers on $\{x,s\}$ (whose colour depends not only on $i\in\{1,2\}$ but also on the parity of $x$ and on the sign of $n$) and may affect the weight which is assigned by the weight function, taking the average of the function $g_{u,x}$ imposes a constraint on the local time at $x$ and the number of $2$-dimers on $\{x,g\}$, while taking the average of the function $g_{x}$ imposes the condition that the local time at $x$ is zero. The domain of the weight function, which has been defined in Section 2, is extended to the negative integers, namely $U:\mathbb{Z}\mapsto\mathbb{R}^{+}_{0}$, under the assumption that $U(n)=0$ for any $n<0$.
Proposition 3.1 (From spins to double dimers).
For $i\in\{1,2\}$, let $u^{i}=(u^{i}_{x})_{x\in V}$, be integer valued vectors and we let $u^{i}_{\pm}=((u^{i}_{\pm})_{x})_{x\in V}$ be such that $(u^{i}_{+})_{x}:=\max\{u^{i}_{x},0\}$ and $(u^{i}_{-})_{x}:=\min\{u^{i}_{x},0\}$. For any $r\in\{1,2\}$, $u\in\mathbb{N}_{0}$, $\beta,h\geq 0$, $A\subset V$, the following identities hold,
$\begin{split}\langle\prod_{x\in V^{e}}e^{i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2})}&\prod_{x\in V^{o}}e^{-i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2})}\rangle_{G,\beta,h,r}\\ =&\frac{1}{\mathcal{Z}_{G,\beta,h,r}}\mu_{\beta,h,r}\Big{(}\mathbbm{1}_{\{\partial w=(u^{1}_{+}-u^{2}_{-},-u^{1}_{-}+u^{2}_{+})\}}\prod_{x\in V}\frac{U\big{(}k_{x}(w)+(u_{-}^{1})_{x}+(u_{-}^{2})_{x}\big{)}}{U\big{(}k_{x}(w)\big{)}}\Big{)},\end{split}$ (3.9)
$\langle$ $\prod_{z\in A}g_{u,z}(s_{z})\rangle_{G,\beta,h,r}=\mathcal{P}_{G,\beta,h,r}\big{(}\forall z\in A,k_{z}(w)=u\big{)},$ (3.10)
$\langle$ $\prod_{z\in A}g_{z}(s_{z})\rangle_{G,\beta,h,r}=\mathcal{P}_{G,\beta,h,r}\big{(}\forall z\in A,n_{z}(w)=0\big{)}.$ (3.11)
In particular we have that $Z_{G,\beta,h,r}=\mathcal{Z}_{G,\beta,h,r}$.
The proof of the proposition is postponed to Section 8. Note that the weight function in the denominator of (3.9) equals zero only for configurations $w$ such that $\mu_{\beta,h,r}(w)=0$. These do not give any contribution to the expectation, hence (3.9) is well defined.
We say that the function $f:\Omega_{s}\mapsto\mathbb{C}$ is good if it admits a Fourier series with real coefficients and, in addition, $\int_{S^{V}}\boldsymbol{ds}|f(\boldsymbol{s})|<\infty.$ For example, for $A\subset V$, the functions $\prod_{x\in A}\sin(s_{x}^{1})$, $\prod_{x\in A}\cos(s_{x}^{1})$ are good. We say that a function $f:\Omega_{s}\mapsto\mathbb{C}$ is positive good if it is good and, in addition, it admits a Fourier series with non negative coefficients. Applying Proposition 3.1 and our assumptions on $U$ we see that if $f$ is good then $\langle f\rangle_{G,\beta,h,r}$ has a real (finite) value and, if $f$ is positive good, it also satisfies $\langle f\rangle_{G,\beta,h,r}\geq 0.$
We now use the double dimer formulation to show that the average of certain observables for the spin system is zero when $h=0$.
Lemma 3.2.
Suppose that $h=0$. For any $\beta>0$, $x,y\in V$, $n\in\mathbb{Z}\setminus\{0\}$ we have that
$\langle e^{i\,n\,s^{i}_{x}}\rangle_{G,\beta,h,r}$ $=0,$ (3.12)
$\langle e^{ni\,(s^{i}_{x}+s^{i}_{y})}\rangle_{G,\beta,h,r}$ $=0.$ (3.13)
Proof.
By Proposition 3.1 we observe that $\langle e^{i\,n\,s^{i}_{x}}\rangle$ is a sum over configurations $(m,\pi)\in\Sigma$ such that $\partial m=(n\delta_{x},0)$ or $\partial m=(0,n\delta_{x})$. This condition implies the existence of a path in the extended graph starting with a dimer on the edge $\{x,s\}$ and ending with a dimer on the edge $\{y,g\}$ for some $y\in V$ (indeed, being the graph bipartite, no path starting on $\{x,s\}$ can end with a dimer of the same colour on the same edge). However, any configuration with a strictly positive number of dimers on some edge $\{y,g\}$ gets weight zero, since $h=0$. This concludes the proof of the first identity. We now prove the second identity. Note that by Proposition 3.1 $\langle e^{\pm ni\,(s^{i}_{x}+s^{i}_{y})}\rangle_{G,\beta,h,r}$ is a sum over configurations $(m,\pi)\in\Sigma$ such that if $x$ and $y$ have the same parity then $\partial m=(n\delta_{x}+n\delta_{y},0)$ or $\partial m=(0,n\delta_{y}+n\delta_{x})$, while if $x$ and $y$ have different parity then $\partial m=(n\delta_{x},n\delta_{y})$ or $\partial m=(n\delta_{y},n\delta_{x})$. Being $h=0$, any such configuration gets weight zero. Indeed, being the graph bipartite, no path can end with dimers on the edges $\{x,s\}$ and $\{y,s\}$ if such dimers have the same colour and $x$ and $y$ have the same parity or if such dimers have different colour and $x$ and $y$ have opposite parity. This concludes the proof. ∎
The next proposition relates the two point functions (3.5), which are defined for the generalized double dimer model, to the two-point functions of the spin system appearing in Theorem 2.1.
We say that the weight function $U$ is nice if $U(0)=1$ and $U(n+1)\leq U(n)$ for each $n\in\mathbb{N}_{0}$. We say that the weight function is fast decaying if it is nice and in addition $U(n+1)\leq K\frac{U(n)}{n}$ for each $n\in\mathbb{N}$, where $K_{U}>0$ is an absolute constant. For example, the weight function of the monomer double dimer model is fast decaying, while the weight function of the double dimer model is not even nice.
Proposition 3.3 (Two-point functions).
For any $x,y\in V$ we have that
[…]
where the first inequality holds for any weight function $U$ (satisfying the general assumptions which have been made in Section 2) if $x$ and $y$ have different parity and for any nice weight function $U$ if $x$ and $y$ have the same parity, the second inequality holds only if $U$ is fast decaying.
Proof.
We omit the subscripts from $\langle\,\,\,\rangle_{G,\beta,0,0}$ for a lighter notation. First note that by Lemma 3.2 we have that
$\langle\cos(\ell s_{x}^{1})\cos(\ell s_{y}^{1})\rangle=\frac{1}{4}\big{(}\langle e^{i\ell(s_{x}^{1}-s_{y}^{1})}\rangle+\langle e^{-i\ell(s_{x}^{1}-s_{y}^{1})}\rangle\big{)}.$
Now if $x$ and $y$ have different parity we obtain from Proposition 3.1 and from the previous expression that
$\langle\cos(\ell s_{x}^{1})\cos(\ell s_{y}^{1})\rangle=\frac{1}{4Z_{G,\beta,0,0}}\Big{(}\sum\limits_{\begin{subarray}{c}(m,\pi)\in\mathcal{W}:\\ \partial m=(\ell\delta_{x}+\ell\delta_{y},0)\end{subarray}}\tilde{\mu}_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m))}{n_{z}(m)!}\\ +\sum\limits_{\begin{subarray}{c}(m,\pi)\in\mathcal{W}:\\ \partial m=(0,\ell\delta_{x}+\ell\delta_{y})\end{subarray}}\tilde{\mu}_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m)-\ell\delta_{x}(z)-\ell\delta_{y}(z))}{n_{z}(m)!}\Big{)}.$ (3.16)
Instead, if $x$ and $y$ have the same parity we obtain from Proposition 3.1 that
$\langle\cos(\ell s_{x}^{1})\cos(\ell s_{y}^{1})\rangle=\frac{1}{4Z_{G,\beta,0,0}}\Big{(}\sum\limits_{\begin{subarray}{c}(m,\pi)\in\mathcal{W}:\\ \partial m=(\ell\delta_{x},\ell\delta_{y})\end{subarray}}\tilde{\mu}_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m)-\ell\delta_{x}(z))}{n_{z}(m)!}\\ +\sum\limits_{\begin{subarray}{c}(m,\pi)\in\mathcal{W}:\\ \partial m=(\ell\delta_{y},\ell\delta_{x})\end{subarray}}\tilde{\mu}_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m)-\ell\delta_{y}(z))}{n_{z}(m)!}\Big{)}.$ (3.17)
If $\ell=1$ we then realise that the first term in the brackets in (3.16) equals $\mu_{G,\beta,0,0}(x\rightarrow y)$. Indeed, any configuration $w=(m,\pi)$ with no dimer touching the ghost vertex and such that $\partial m=(\delta_{x}+\delta_{y},0)$ (with a $1$-dimer on $\{x,s\}$ and on $\{y,s\}$) necessarily has a path from $\{x,s\}$ to $\{y,s\}$ if $x$ and $y$ have different parity. Similarly, using the monotonicity property of $U$, we see that the first term in the brackets in (3.17) is greater or equal than $\mu_{G,\beta,0,0}(x\rightarrow y)$. Indeed, any configuration $w=(m,\pi)$ with no dimer touching the ghost vertex and such that $\partial m=(\delta_{x},\delta_{y})$ (with a $1$-dimer on $\{x,s\}$ and a $2$-dimer on $\{y,s\}$) necessarily has a path from $\{x,s\}$ to $\{y,s\}$ if $x$ and $y$ have the same parity. In both cases, being the second term in the brackets non negative, we conclude the proof of (3.14).
[…]
  1. (i)
    switching to blue (resp. red) the colour of each red (resp. blue) dimer belonging to one of the two open paths from $\{x,s\}$ to $\{y,s\}$,
  2. (ii)
    removing the two dimers on $\{x,s\}$ and the two dimers on $\{y,s\}$,
  3. (iii)
    for each $z\in\{x,y\}$ pairing together at $z$ the two dimers which were paired to those which have been removed.
Thus, $f(\mathcal{A})=\{x\leftrightarrow y\}\subset\mathcal{W}^{\ell}$ and $f^{-1}(\{x\leftrightarrow y\})=\mathcal{A}$. We observe that for any configuration $w=(m,\pi)\in\{x\leftrightarrow y\}\subset\mathcal{W}^{\ell}$ and $w^{\prime}=(m^{\prime},\pi^{\prime})\in f^{-1}(w)$ we have that, for each $z\in V$,
$|f^{-1}(w)|\geq 1,\quad\quad n_{z}(m^{\prime})=n_{z}(m)+\delta_{x}(z)+\delta_{y}(z)\quad\quad\tilde{\mu}_{\beta,h,r}(m)=\mu_{\beta,h,r}(m^{\prime}).$ (3.18)
Using (3.18) for the first inequality and the property that $U$ is fast decaying for the second inequality we obtain that
$\mathcal{G}^{(2)}_{G,\beta}(x,y)=$ $\sum\limits_{w=(m,\pi)\in\{x\leftrightarrow y\}}\tilde{\mu}_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m))}{n_{z}(m)!}$
$=$ $\sum\limits_{w=(m,\pi)\in\{x\leftrightarrow y\}}\frac{1}{|f^{-1}(w)|}\sum\limits_{w^{\prime}=(m^{\prime},\pi^{\prime})\in f^{-1}(w)}\tilde{\mu}_{\beta,0,0}(m)\prod_{z\in V}\frac{U\big{(}n_{z}(m)\big{)}}{n_{z}(m)!}$
$\leq$ $\sum\limits_{w=(m,\pi)\in\{x\leftrightarrow y\}}\sum\limits_{w^{\prime}=(m^{\prime},\pi^{\prime})\in f^{-1}(w)}\tilde{\mu}_{\beta,0,0}(m^{\prime})\prod_{z\in V}\frac{U\big{(}n_{z}(m^{\prime})-\delta_{x}(z)-\delta_{y}(z)\big{)}}{\big{(}n_{z}(m^{\prime})-\delta_{x}(z)-\delta_{y}(z)\big{)}!}$
$\leq$ $K_{U}\,\,\sum\limits_{w^{\prime}=(m^{\prime},\pi^{\prime})\in\mathcal{A}}\tilde{\mu}_{\beta,0,0}(m^{\prime})\prod_{z\in V}\frac{U\big{(}n_{z}(m^{\prime})-2\delta_{x}(z)-2\delta_{y}(z)\big{)}}{n_{z}(m^{\prime})!},$
recognizing that the last term corresponds to the second term in the brackets in (3.16) and then concluding the proof of (3.15).
4 Reflection positivity and its consequences
In this section we show that the complex measure $\langle\,\,\,\rangle_{L,K,\beta,h,r}$ which has been introduced in Section 2 is reflection positive and use this property to derive convexity and monotonicity properties for the two-point function, following [23], and to derive a Chessboard estimate, which is a well known consequence. The reflection positivity property of the complex measure reflects an equivalent reflection positivity property of the double dimer model [30]. It relies on the complex duality relation between vertices having opposite parity in (2.4) and its proof is classical [32, Chapter 10].
In the whole section we fix arbitrary values $\beta,h,r\geq 0$, $L,K\in 2\mathbb{N}\cup\{1\}$, and sometimes omit the subscripts of $\langle\,\,\,\rangle_{L,K,\beta,h,r}$ for a lighter notation.
4.1 Reflection positivity
Recall that $\mathbb{S}_{L,K}$ is the slab torus. We let $\mathbb{S}_{L,K}^{o}\subset\mathbb{S}_{L,K}$ be the set of odd sites, and $\mathbb{S}_{L,K}^{e}\subset\mathbb{S}_{L,K}$ be the set of even sites.
We let $\mathcal{S}$ be a plane which is orthogonal to the Cartesian vector $e_{i}$ for some $i\in\{1,2,3\}$ and which intersects the midpoint of some edges. We let $\Theta:\mathbb{S}_{L,K}\mapsto\mathbb{S}_{L,K}$ be a reflection operator with respect to such plane.
Let us identify two (disjoint, connected) torus halves, $\mathbb{S}_{L,K}^{\pm}$, such that $\Theta(\mathbb{S}_{L,K}^{\pm})=\mathbb{S}_{L,K}^{\mp}$. By a slight abuse of notation we keep using $\Theta:\Omega_{s}\mapsto\Omega_{s}$ to denote a reflection operator that associates to each configuration $\boldsymbol{s}=(s_{x})_{x\in V}$ the reflected spin configuration $\Theta(\boldsymbol{s})=(s_{\Theta(x)})_{x\in\mathbb{S}_{L,K}}$.
Given a function $f:\Omega_{s}\mapsto\mathbb{C}$, we define the reflected function $\Theta f$ as,
$\forall\boldsymbol{s}\in\Omega_{s}\quad\Theta f(\boldsymbol{s}):={f(\Theta(\boldsymbol{s}))}.$
We say that the function $f:\Omega_{s}\mapsto\mathbb{C}$ has domain $A\subset\mathbb{S}_{L,K}$ if for each pair of configurations $\boldsymbol{s},\boldsymbol{s}^{\prime}$ such that $s_{x}=s_{x}^{\prime}$ for each $x\in A$, $f(\boldsymbol{s})=f(\boldsymbol{s}^{\prime})$. Let $\mathcal{A}^{\pm}$ be the set of functions $f:\Omega_{s}\mapsto\mathbb{C}$ which have domain $\mathbb{S}_{L,K}^{\pm}$,.
Proposition 4.1.
For any $f,g\in\mathcal{A}^{+}$ we have
  1. (i)
    $\langle f\,\,\overline{\Theta g}\rangle_{L,K,\beta,h,r}={\langle\overline{g}\,\,\Theta f\rangle}_{L,K,\beta,h,r},$
  2. (ii)
    $\langle f\,\overline{\Theta f}\rangle_{L,K,\beta,h,r}\geq 0$.
This in turn implies that
${\Big{(}{{Re}\,\langle f\,\overline{\Theta g}\,\rangle}_{L,K,\beta,h,r}\Big{)}}^{2}\leq\langle\,f\,\overline{\Theta f}\rangle_{L,K,\beta,h,r}\,\langle g\,\overline{\Theta g}\rangle_{L,K,\beta,h,r}.$ (4.1)
Proof.
This proof strategy is classical and can be found in several papers about reflection positivity, see for example [32, Chapter 10]. We observe that the property (i) simply follows by symmetry, hence we turn to the proof of (ii). Let us define for any measurable function $f:\Omega_{s}\mapsto\mathbb{C}$ the operator
[…]
and define $\Omega_{\pm}=S^{\mathbb{S}^{\pm}_{L,K}}$. By symmetry we have that
[…]
Now let $E^{\pm}$ be the subset of of edges of $\mathbb{S}_{L,K}$ which are entirely contained in $\mathbb{S}_{L,K}^{\pm}$ and $E_{R}$ be the set of edges intersecting both $\mathbb{S}_{L,K}^{-}$ and $\mathbb{S}_{L,K}^{+}$. We observe that we can write
$\langle f\,\,\overline{\Theta f}\rangle_{L,K,\beta,h,r}=\langle f\,\,\overline{\Theta f}\,\,e^{\mathcal{H}_{+}}e^{\mathcal{H}_{-}}e^{\sum_{\{x,y\}\in E_{R}}\beta e^{is_{x}^{1}}\,\,\Theta(e^{is_{x}^{1}})}e^{\sum_{\{x,y\}\in E_{R}}\beta e^{-is_{x}^{2}}\,\,\Theta(e^{-is_{x}^{2}})}\rangle_{0}$ (4.3)
where
±$\mathcal{H}_{\pm}(\boldsymbol{s}):=\sum\limits_{\{i,j\}\in E^{\pm}}e^{i(s_{i}^{1}-s_{j}^{1})}+\sum\limits_{\{i,j\}\in E^{\pm}}e^{-i(s_{i}^{2}-s_{j}^{2})}+\,h\sum\limits_{k\in\mathbb{S}_{L,K}^{\pm}}\,2\,\cos(rs^{1}_{k})$
and we used the convention that if $\{x,y\}\in E_{R}$, then $x\in\mathbb{S}_{L,K}^{+}$ and $y\in\mathbb{S}_{L,K}^{-}$. Noting that $e^{\mathcal{H}_{\pm}}=\overline{\Theta e^{\mathcal{H}_{\mp}}}$ and expanding the last two exponentials in (4.3) we obtain that
$\langle f\,\,\overline{\Theta f}\rangle_{L,K,\beta,h,r}\\ =\sum\limits_{m^{1},m^{2}\in{\mathbb{N}_{0}}^{E_{R}}}\Big{\langle}f\,\,\overline{\Theta f}\,\,e^{\mathcal{H}_{+}}\,\Theta\overline{\big{(}e^{\mathcal{H}_{+}}\big{)}}\prod_{\{x,y\}\in E_{R}}\frac{\big{(}\beta e^{is_{x}^{1}}\,\,\Theta(\overline{e^{is_{x}^{1}}})\Big{)}^{m_{\{x,y\}}^{1}}}{m_{\{x,y\}}^{1}!}\frac{\big{(}\beta e^{-is_{x}^{2}}\,\,\Theta(\overline{e^{-is_{x}^{2}}})\Big{)}^{m_{\{x,y\}}^{2}}}{m^{2}_{\{x,y\}}!}\Big{\rangle}_{0}.$ (4.4)
For each term in the previous sum we set $h=fe^{\mathcal{H}_{+}}\prod_{\{x,y\}\in E_{R}}e^{im^{1}_{\{x,y\}}s_{x}^{1}}e^{-im^{2}_{\{x,y\}}s_{x}^{2}}.$ We then note that $h\in\mathcal{A}^{+}$ and that $\langle f\,\,\overline{\Theta f}\rangle_{L,K,\beta,h,r}$ can be expressed as a sum of terms of the form $\langle h\,\,\overline{\Theta(h)}\rangle_{0}$ multiplied by positive constants. By applying (4.2) for each term in the previous sum we then deduce (ii), concluding the proof. ∎
4.2 Convexity and monotonicity of two-point functions
The vector of functions $F=(F_{x})_{x\in\mathbb{S}_{L,K}}$, with $F_{x}:S\mapsto\mathbb{R}$, is said to be reflection invariant if for any $x,y,z\in\mathbb{S}_{L,K}$ the two following conditions hold
  1. (i)
    $\Theta F_{x}=F_{\Theta x}$ for any reflection through edges $\Theta$,
  2. (ii)
    $\langle F_{x}F_{y}\rangle=\langle F_{x+z}F_{y+z}\rangle\geq 0$, where the sum is with respect to the torus metric.
The next lemma will allow us to derive convexity properties for two-point functions, such us the functions in the RHS of (3.14) and (3.15).
Lemma 4.2.
Consider any reflection plane $\mathcal{S}$, let $\Theta$ be the corresponding reflection operator and $\mathbb{S}_{L,K}^{\pm}\subset\mathbb{S}_{L,K}$ the corresponding torus halves as in Section 4.1. Let $(F_{x})_{x\in\mathbb{S}_{L,K}}$ be a reflection invariant vector of functions such that $F_{o}$ has domain $D\subset\mathbb{S}_{L,K}$, with $o\in D$, and define $\ell:=\sup\{d(o,x):x\in D\}$. Let $Q\subset\mathbb{S}_{L,K}$, $Q^{\pm}:=(Q\cap\mathbb{S}_{L,K}^{\pm})\cup\Theta(Q\cap\mathbb{S}_{L,K}^{\pm})$, assume that $|Q^{+}|=|Q^{-}|$ and that $d_{H}(Q^{\pm},\mathcal{S})>\ell$, where $d_{H}$ is the Hausdorff distance with respect to the L2 norm. Then,
$2\sum\limits_{\begin{subarray}{c}x,y\in Q:\\ x\neq y\end{subarray}}\langle F_{x}F_{y}\rangle_{L,K,\beta,h,r}\leq\sum\limits_{\begin{subarray}{c}x,y\in Q^{+}:\\ x\neq y\end{subarray}}\langle F_{x}F_{y}\rangle_{L,K,\beta,h,r}+\sum\limits_{\begin{subarray}{c}x,y\in Q^{-}:\\ x\neq y\end{subarray}}\langle F_{x}F_{y}\rangle_{L,K,\beta,h,r}.$ (4.11)
Proof.
Let $\varphi>0$ be an arbitrary real number. From Proposition 4.1 we deduce that,
$\langle\prod_{x\in Q}(1+\varphi F_{x})\rangle^{2}\leq\langle\prod_{x\in Q^{+}}(1+\varphi F_{x})\rangle\langle\prod_{x\in Q^{-}}(1+\varphi F_{x})\rangle.$
Since the inequality holds for each $\varphi$, we perform an expansion in the limit as $\varphi\rightarrow 0$ and we obtain (4.11) from a comparison of the terms $O(\varphi^{2})$. ∎
The next proposition derives a convexity property along the Cartesian axis for two point functions. Our result holds for two-point functions which can be expressed as the expectation of the product of two functions belonging to a reflection invariant vector.
Proposition 4.3 (Convexity).
Let $F=(F_{x})_{x\in\mathbb{S}_{L,K}}$ be a reflection invariant vector of functions satisfying the assumptions in Lemma 4.2. Define for any $x\in\mathbb{S}_{L,K}$
$\mathcal{G}(x):=\langle F_{o}F_{x}\rangle_{L,K,\beta,h,r}$ (4.12)
Let $x_{n}$ correspond to the vertex $(n,0,0)$ or to the vertex $(0,n,0)$, let $y_{k}$ correspond to the vertex $(0,0,k)$. For odd vertices $n,k$ such that $\min\{2\ell+1,3\}\leq n\leq L-\min\{2\ell+1,3\}$ and $\min\{2\ell+1,3\}\leq k\leq K-\min\{2\ell+1,3\}$ we have that,
$\begin{split}2\,\mathcal{G}(x_{n})&\leq\mathcal{G}(x_{n+2})+\mathcal{G}(x_{n-2}),\\ 2\,\mathcal{G}(y_{k})&\leq\mathcal{G}(y_{k+2})+\mathcal{G}(y_{k-2}).\end{split}$ (4.13)
Proof.
The proposition follows from am immediate application of Lemma 4.2 with $\mathcal{S}=\{(x_{1},x_{2},x_{3})\in\mathbb{S}_{L,K}\,:\,x_{1}=\frac{n}{2}-1\}$ and $Q=\{(0,0,0),(n,0,0)\}$, or $\mathcal{S}=\{(x_{1},x_{2},x_{3})\in\mathbb{S}_{L,K}\,:\,x_{2}=\frac{n}{2}-1\}$ and $Q=\{(0,0,0),(0,n,0)\}$ or $\mathcal{S}=\{(x_{1},x_{2},x_{3})\in\mathbb{S}_{L,K}\,:\,x_{3}=\frac{k}{2}-1\}$ and $Q=\{(0,0,0),(0,0,k)\}$. ∎
The next property is a consequence of convexity, namely it is a monotonicity property for the two-point function.
Proposition 4.4 (Monotonicity).
Under the same assumptions as in Proposition 4.3 we have that, for any odd integer $n\in[0,\frac{L}{2}-1)$ and $k\in[0,\frac{K}{2}-1)$,
$\mathcal{G}(x_{n})\geq\mathcal{G}(x_{n+2})\quad\quad\mathcal{G}(x_{k})\geq\mathcal{G}(x_{k+2}).$
For the proof of the proposition (given the analogous of Proposition 4.1) we refer to [23] or to [24, Theorem 2.1].
4.3 Chessboard estimates
A classical consequence of reflection positivity is the chessboard estimate. Let $f$ be a function having domain $\{o\}$, let $e_{1},e_{2}\ldots e_{k}$ a sequence of edges of the graph $\mathbb{S}_{L,K}$ forming a self-avoiding path from $o$ to $t\in\mathbb{S}_{L,K}$, let $\Theta_{1},\ldots\Theta_{k}$ be a sequence of reflection through edges such that $\Theta_{i}$ is a reflection with respect to the plane orthogonal to $e_{i}$ and intersecting the midpoint of such an edge for each $i\in\{1,\ldots,k\}$. Define,
$f^{t}:=\begin{cases}\overline{\Theta_{k}\circ\Theta_{k-1}\,\ldots\,\circ\Theta_{1}\,(f)}&\mbox{ if $k$ is even,}\\ \Theta_{k}\circ\Theta_{k-1}\,\ldots\,\circ\Theta_{1}\,(f)&\mbox{ if $k$ is odd.}\end{cases}$
Observe that the function $f^{t}$ does not depend on the chosen path and has domain $t$.
Proposition 4.5 (Chessboard estimate).
Let $f=(f_{t})_{t\in\mathbb{S}_{L,K}}$ be a sequence of good functions having domain $\{o\}$. We have that,
$\langle\prod_{t\in\mathbb{S}_{L,K}}f^{t}_{t}\rangle_{L,K,\beta,h,r}\,\leq\,\prod_{t\in\mathbb{S}_{L,K}}\,\,{\big{\langle}\,\prod_{s\in\mathbb{S}_{L,K}}f_{t}^{s}\,\big{\rangle}_{L,K,\beta,h,r}}^{\frac{1}{|\mathbb{S}_{L,K}|}}$
Given Proposition 4.1, the proof of Propoposition 4.5 is classical and is based on performing a sequence of reflections, we refer for example to [32, Theorem 10.11] for its proof.
5 Probabilistic estimates
In this section we use reflection positivity and probabilistic arguments to provide basic estimates for some quantities of interest which are defined in the framework of the spin system. Most of our proofs however are formulated in the framework of the double dimer model.
In the whole section we use $G=(V,E)$ to denote the graph corresponding to the slab torus $\mathbb{S}_{L,K}$, we let $d$ be the degree of the origin and we omit the subscripts from $\langle\,\,\,\rangle_{L,K,\beta,h,r}$ when unnecessary for a lighter notation.
The following technical lemma is a consequence of the chessboard estimate.
Lemma 5.1 (Chessboard upper bound).
Let $f_{1},f_{2}:[0,2\pi)\mapsto\mathbb{R}$ be two good functions which admit a Fourier series with a bounded numbers of terms. For any $\beta>0$, there exists $c=c(\beta,f_{1},f_{2})\in(0,\infty)$ such that, for any $h\in[0,1]$, $i,j\in\{1,2\}$, $L,K\in 2\mathbb{N}\cup\{1\}$, and disjoint sets $A,B\subset\mathbb{S}_{L,K}$
$\langle\prod_{x\in A}f_{1}(s_{x}^{i})\prod_{x\in B}f_{2}(s_{x}^{j})\rangle_{L,K,\beta,h,r}\leq c^{|A|+|B|}.$ (5.1)
Proof.
From Proposition 4.5 we deduce that
$\langle\prod_{x\in A}f_{1}(s_{x}^{i})\prod_{x\in B}f_{2}(s_{x}^{j})\rangle\leq\Big{(}\big{\langle}\prod_{x\in V}f_{1}(s_{x}^{i})\big{\rangle}\Big{)}^{\frac{|A|}{|V|}}\Big{(}\big{\langle}\prod_{x\in V}f_{2}(s_{x}^{j})\big{\rangle}\Big{)}^{\frac{|B|}{|V|}}$ (5.2)
Using that by assumption for each $k\in\{1,2\}$ we can write $f_{k}(s_{x}^{i})=\sum_{n=-K}^{K}a_{n}e^{i\,n\,s_{x}^{i}}$ some $K<\infty$ and some real coefficients $a_{-K},\ldots a_{K}$ and calling $M=\max\{|a_{n}|\,:-K\leq n\leq K\}$ we obtain from Proposition 3.1 that
$\langle\prod_{x\in V}f_{k}(s^{i}_{x})\rangle_{L,K,\beta,h,r}=\sum\limits_{\boldsymbol{n}\in\{-K,\ldots,K\}^{V}}\langle\prod_{x\in V}a_{n_{x}}e^{i\,\,n_{x}^{i}\,s^{i}_{x}}\rangle\rangle_{L,K,\beta,h,r}$
$\leq{(2K+1)}^{|V|}\,\,{\max\{M,1\}}^{|V|}\max\Big{\{}\frac{1}{Z_{G,\beta,h,r}}\sum\limits_{\begin{subarray}{c}m\in\Sigma:\\ \partial m=(\boldsymbol{n}^{1},\,\,\boldsymbol{n}^{2})\end{subarray}}\tilde{\mu}_{\beta,h,r}(m)\,\,:\,\,\boldsymbol{n}=(\boldsymbol{n^{1}},\boldsymbol{n^{2}})\in\{-K,\ldots,K\}^{V\times V}\Big{\}}$
$\leq c^{|V|}{(2K+1)}^{|V|}\,\,{\max\{M,1\}}^{|V|}e^{2\beta d|V|+2\beta h|V|}.$
For the first inequality we used assumption that $U(n)\leq 1$ for each $n\in\mathbb{N}_{0}$. For the second inequality we summed over the number of dimers independently for each edge $e\in E\cup E_{g}$ (ignoring any constraint on their cardinality) and we used that $Z_{G,\beta,h,r}\geq c^{|V|}$, for some $c=c(\beta)\in(0,\infty)$, concluding the proof. ∎
Lemma 5.2.
For any $\beta>0$, $\epsilon>0$, there exists a constant $c=c(\beta)>0$ and $M=M(\beta,\epsilon)<\infty$ such that for any $h\in[0,1]$, $r\in\{1,2\}$, $L,K\in 2\mathbb{N}\cup\{1\}$, $z\sim o$,
$\mathcal{P}_{L,K,\beta,h,r}\big{(}m_{\{o,z\}}^{i}>M\big{)}$ $\leq\epsilon,$ (5.3)
$\mathcal{P}_{L,K,\beta,h,r}\big{(}n_{o}^{i}>0\big{)}$ $\geq c.$ (5.4)
Proof.
For the first identity below we use Proposition 3.1, for the third inequality we use Proposition 4.5, obtaining that
$\mathcal{P}_{L,K,\beta,h,r}\big{(}m_{\{o,z\}}^{i}>M\big{)}$ $\leq\mathcal{P}_{L,K,\beta,h,r}\big{(}n_{o}-m_{\{o,g\}}^{2}>M\big{)}$
$\leq\sum\limits_{u=M+1}^{\infty}\mathcal{P}_{L,K,\beta,h,r}\big{(}n_{o}-m_{\{o,g\}}^{2}=u\big{)}$
$=\sum\limits_{u=M+1}^{\infty}\langle\,g_{u,o}(s_{o})\,\big{\rangle}$
$\leq\sum\limits_{u=M+1}^{\infty}\Big{(}\langle\prod_{x\in V}g_{u,x}(s_{x})\big{\rangle}\Big{)}^{\frac{1}{|V|}}$
$=\sum\limits_{u=M+1}^{\infty}\mathcal{P}_{L,K,\beta,h,r}\big{(}\forall x\in V:n_{x}^{i}-m_{\{x,g\}}^{2}=u\big{)}^{\frac{1}{|V|}}$
$\leq\sum\limits_{u=M+1}^{\infty}\Big{(}\frac{1}{Z_{L,K,\beta,h,r}}d^{|V|}\frac{\max\{\beta,1\}^{u|V|}\,e^{2(|E|+|E_{g}|)\max\{\beta,1\}}}{\big{(}(u/d)!\big{)}^{\frac{1}{d}}}\Big{)}^{\frac{1}{|V|}}$
$\leq c\sum\limits_{u=M+1}^{\infty}\frac{1}{\big{(}(u/d)!\big{)}^{\frac{1}{d}}}$
For the last inequality we used a crude upper bound for the probability of the event $\{\forall x\in V:n_{x}-m_{\{x,g\}}^{2}=u\}$ by using the fact that, if $n_{x}(w)-m_{\{x,g\}}^{2}(w)=u$, then there exists one of the $d$ edges of $E$ touching $x$ with at least $u/d$ dimers of colour $1$ and there are at most $d$ possibilities for choosing such edge, and for each remaining edge and each colour we summed the number of dimers independently (using $U\leq 1$ and ignoring the constraint on the dimer cardinality) getting a factor $e^{\beta}$. For the last inequality we also used that $Z_{L,K,\beta,h,r}\geq c^{|V|}$ for some $c=c(\beta)$. Being the sum convergent we can make it arbitrarily small choosing $M$ large, obtaining (5.3).
For (5.4) we observe that if $U(0)=0$ then the proof is obvious by definition of the measure since by our assumption there exists $k>0$ such that $U(k)>0$. Hence, we can assume that $U(0)>0$. By Proposition 3.1 and Proposition 4.5 we obtain that
$\mathcal{P}_{L,K,\beta,h,r}\big{(}n_{o}=0\big{)}$ $=\langle g_{o}(s_{o})\big{\rangle}$
$\leq\Big{(}\langle\prod_{x\in V}g_{x}(s_{x})\big{\rangle}\Big{)}^{\frac{1}{|V}|}$
$=\Big{(}\mathcal{P}_{L,K,\beta,h,r}\big{(}\forall x\in V,n_{x}=0\big{)}\Big{)}^{\frac{1}{|V|}}$
$=\Big{(}\frac{U(0)^{|V|}}{Z_{G,\beta,h,r}}\Big{)}^{\frac{1}{|V|}}<c<1.$
To show that $c<1$ we identify a subset $A\subset V$ such that $|A|>\frac{1}{100}|V|$ and $d(x,y)>2$ for each distinct $x,y\in A$. We then lower bound $Z_{G,\beta,h,r}$ by considering only the contribution of the configurations which have the vertices of $A$ either empty or touched by precisely $k$ blue and $k$ red dimers on the same edge and all the remaining vertices empty, hence obtaining $Z_{G,\beta,h,r}\geq U(0)^{|V|}(1+\frac{\beta^{2k}U(k)^{2}}{k!^{2}U(0)^{2}})^{{\frac{|V|}{100}}}$. This leads to the desired bound and concludes the proof. ∎
Lemma 5.3.
For any $\beta>0$, $h\in[0,1]$, $r\in\mathbb{N}$, there exists a constant $c=c(\beta)\in(0,\infty)$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$, for any $z\sim o$, we have that,
$\langle e^{i(s^{i}_{o}-s^{i}_{z})}\rangle_{L,K,\beta,h,r}\,\,\,$ $=\frac{1}{\beta}\mathcal{E}_{\mathbb{S}_{L,K},\beta,h,r}\big{(}\frac{1}{m_{\{o,z\}}^{i}}\mathbbm{1}_{\{m_{\{o,z\}}^{i}>0\}}\big{)}\leq\frac{1}{\beta},$ (5.5)
$\sum\limits_{z\sim o}\langle e^{i(s^{i}_{o}-s^{i}_{z})}\rangle$ $\geq c.$ (5.6)
Proof.
We start from the proof of the identity in the statement. From Proposition 3.1 we deduce that
$\langle e^{i(s^{i}_{o}-s^{i}_{z})}\rangle_{L,K,\beta,h,r}=\frac{\mu_{\beta,h,r}(\mathcal{A})}{\mathcal{Z}_{G,\beta,h,r}},$
where $\mathcal{A}$ is the set of configurations with precisely one $i$-dimer on $\{o,s\}$, one $i$-dimer on $\{z,s\}$, and no further dimer touching $s$.
We introduce a map $f$ which takes any configuration $(m,\pi)\in\mathcal{A}$ and associates to it the configuration $f(m,\pi)\in\mathcal{W}^{\ell}$ which is defined by following the next three steps:
  1. (i)
    removing the $i$-dimers on $\{o,s\}$ and $\{z,s\}$,
  2. (ii)
    adding a $i$-dimer on $\{o,z\}$ and assigning to it the label $m^{i}_{\{o,z\}}+1$,
  3. (iii)
    pairing the added dimer at $z$ and at $o$ to the dimers which were paired to those that have been removed.
It follows from the definition that the map is a bijection from $\mathcal{A}$ to $f(\mathcal{A})=\{(m,\pi)\in\mathcal{W}^{\ell}\,:\,m^{1}_{\{o,z\}}>0\}.$ For any configuration $w=(m,\pi)\in\mathcal{A}$, $w^{\prime}=(m^{\prime},\pi^{\prime})=f(w)$, we have that
$\mu_{\beta,h,r}(w)=\frac{1}{\beta}\frac{1}{(m^{i}_{\{o,z\}})^{\prime}}\mu_{\beta,h,r}(w^{\prime})$ (5.7)
From this we then deduce the identity (5.5). The upper bound in (5.5) then follows trivially.
Let us now prove (5.6). From Lemma 5.2 and (5.5) we deduce that
$\sum\limits_{z\sim o}\langle e^{i(s^{i}_{o}-s^{i}_{z})}\rangle=\sum\limits_{z\sim o}\frac{1}{\beta}\mathcal{E}_{L,K,\beta,h,r}\big{(}\frac{1}{m_{\{o,z\}}^{i}}\mathbbm{1}_{\{m_{\{o,z\}}^{i}>0\}}\big{)}\geq\sum\limits_{z\sim o}\frac{1}{M\beta}\mathcal{P}_{L,K,\beta,h,r}\big{(}0<m_{\{o,z\}}^{i},n_{o}^{i}\leq M\big{)}\\ \geq\sum\limits_{z\sim o}\frac{1}{M\beta}\Big{(}\mathcal{P}_{L,K,\beta,h,r}\big{(}m_{\{o,z\}}^{i}>0\big{)}-\mathcal{P}_{L,K,\beta,h,r}\big{(}n_{o}^{i}>M\big{)}\Big{)}\\ \geq\frac{1}{M\beta}\Big{(}\mathcal{P}_{L,K,\beta,h,r}\big{(}n_{o}^{i}>0\big{)}-6\mathcal{P}_{L,K,\beta,h,r}\big{(}n_{o}^{i}>M\big{)}\Big{)}\geq c.$
where the last inequality follows from Lemma 5.2 by choosing $M$ large enough depending on $\beta$.
We now introduce an important quantity of interest, the magnetisation,
$m_{G}(\beta,h,r):=\,\langle\cos(rs_{o}^{1})\rangle_{G,\beta,h,r}.$
In the next proposition we explore the limiting behaviour $|\mathbb{S}_{L,K}|\rightarrow\infty$ and $h\sim\frac{1}{|\mathbb{S}_{L,K}|}\rightarrow 0$ of this quantity. In the first claim of the proposition we show its important connection to the two point functions which have been analysed in Proposition 3.3 in the case $h=0$. Thanks to (5.8) below we will deduce an upper bound on the Cesaro sum of two point functions by bounding the magnetisation from above .
Proposition 5.4 (Magnetisation expansion).
Set $\beta\geq 0$, $r\in\{1,2\}$, $\ell\in\mathbb{N}$. There exist a constant $c=c(\beta,\ell)\in(0,\infty)$ and a value $h_{0}>0$ small enough such that for any sufficiently small $\tilde{h}\in(0,h_{0})$, every $L,K\in 2\mathbb{N}\cup\{1\}$
$m_{\mathbb{S}_{L,K}}(\beta,h_{L,K},r)$ $\geq\,\,\frac{\beta\tilde{h}}{|\mathbb{S}_{L,K}|}\sum\limits_{x\in\mathbb{S}_{L,K}}\langle\cos(rs_{o}^{1})\cos(rs_{x}^{1})\rangle_{L,K,\beta,0,0},$ (5.8)
$\langle\cos(\ell s_{o}^{1})\rangle_{L,K,\beta,h_{L,K},r}$ $\leq c\,\tilde{h}$ (5.9)
where $h_{L,K}=\frac{\tilde{h}}{|\mathbb{S}_{L,K}|}$.
Proof.
From a Taylor expansion of the exponentials in (2.4) and Lemma 3.2 we deduce that
$\langle\cos(\ell\,s_{o}^{1})\rangle_{L,K,\beta,h_{L,K},r}=\sum\limits_{\begin{subarray}{c}m\in\mathbb{N}_{0}^{V}:\\ m\neq 0\end{subarray}}\langle\cos(\ell s_{o}^{1})\prod_{x\in V}\frac{\big{(}2\beta h_{L,K}\cos(rs_{x}^{1})\big{)}^{m_{x}}}{m_{x}!}\rangle_{L,K,\beta,0,r}.$ (5.10)
For a lower bound we use the positive goodness condition (introduced right after Proposition 3.1) and reduce the sum in (5.10) to the terms such that $m_{x}=1$ for just one $x\in V$ and $m_{y}=0$ for all $y\neq x$, thus obtaining (5.8) using the definition of $h_{L,K}$ and that $\ell=r$. For (5.9) we start from (5.10) and use Lemma 5.1 to deduce that
$\langle\cos(\ell s_{o}^{1})\rangle_{L,K,\beta,h_{L,K},r}\leq c\sum\limits_{\begin{subarray}{c}m\in\mathbb{N}_{0}^{V}:\\ m\neq 0\end{subarray}}\prod_{x\in V}\frac{\big{(}\beta h_{L,K}c\big{)}^{m_{x}}}{m_{x}!}=c(e^{c\,|V|\,\beta\,h_{L,K}}-1)\leq 2\,\beta\,c^{2}\tilde{h},$
where in the last step we used that $\tilde{h}$ is small enough depending on $\beta$. ∎
6 Cauchy-Schwarz inequality
This section presents the central argument for our new proof of the Merming Wagner theorem for non positive (but reflection positive) measures.
We denote by ${\mathcal{H}}:=(-L/2,L/2]\cap(2\mathbb{Z}+1)e_{1}$ and by ${\mathcal{V}}:=(-L/2,L/2]\cap(2\mathbb{Z}+1)e_{2}$, $\mathcal{W}:=(-\frac{K}{2},\frac{K}{2}]\cap(2\mathbb{Z}+1)e_{3}$ the subsets of $\mathbb{S}_{L,K}$ along the line passing through the origin and parallel to the Cartesian vectors. We introduce the Fourier dual torus,
$\mathbb{S}_{L,K}^{*}:=\big{\{}2\pi(\frac{n_{1}}{L},\frac{n_{2}}{L},\frac{n_{3}}{K})\in\mathbb{R}^{3}\,\,:\,\,n_{i}\in(-\frac{L}{2},\frac{L}{2}]\cap\mathbb{Z}\,,\text{ for }i=1,2\,,n_{3}\in(-\frac{K}{2},\frac{K}{2}]\cap\mathbb{Z}\big{\}},$
and, given and fixed any Fourier mode $k\in\mathbb{S}_{L,K}^{*}$, for any two vectors $a,b\in\mathbb{C}^{\mathbb{S}_{L,K}}$, we introduce the inner product,
$a\circ b:=\sum\limits_{x\in\mathcal{H}\cup\mathcal{V}\cup\mathcal{W}}\cos(k\cdot x)\,a_{o}\,b_{x}+a_{o}b_{o}$
We now define the extended Hamiltonian,
$\tilde{H}:=H+\sum\limits_{x\in\mathbb{S}^{e}_{L,k}}\log(g_{x})+\sum\limits_{x\in\mathbb{S}^{o}_{L,k}}\log(\overline{g_{x}}),$
and the vector functions,
$A:=(\sin(rs_{x}^{1}))_{x\in\mathbb{S}_{L,K}}\quad\quad B:=\beta\,\big{(}\partial_{s_{x}^{1}}\tilde{H}-\partial_{s_{x}^{2}}\tilde{H}\big{)}_{x\in\mathbb{S}_{L,K}},$
and the quantity
$p_{z}=p_{z}(L,K,\beta,h,r):=\beta\big{(}\langle e^{i(s_{o}^{1}-s_{z}^{1})}\rangle_{L,K,\beta,h,r}+\langle e^{-i(s_{o}^{2}-s_{z}^{2}}\rangle_{L,K,\beta,h,r}\big{)},$
for any $z\sim o$. Standard computations based on integration by parts lead to the following central identities, which are one of the few results of our paper which rely in an essential way on the spin representation.
Lemma 6.1 (Identities).
For any $\beta,h\geq 0$, $r\in\{1,2\}$, $L,K\in 2\mathbb{N}\cup\{1\}$, we have that,
$\langle A\circ B\rangle_{L,K,\beta,h,r}$ $=\langle B\circ A\rangle_{L,K,\beta,h,r}=r\,m_{\mathbb{S}_{L,K}}(\beta,h,r)$ (6.1)
$\langle B\circ B\rangle_{L,K,\beta,h,r}$ $=\sum\limits_{z\sim o}p_{z}\,\big{(}1-\cos(k\cdot z)\big{)}+2r^{2}\,h\,m_{\mathbb{S}_{L,K}}(\beta,h,r)$ (6.2)
Proof.
[…]
$\langle\sin(rs_{o}^{1})\beta\,\big{(}\partial_{s_{x}^{1}}\tilde{H}-\partial_{s_{x}^{2}}\tilde{H}\big{)}\rangle=r\langle\cos(rs_{o}^{1})\rangle\,\,\delta_{o}(x).$
For the second identity we use again integration by parts and the periodicity of the function in the average to conclude that,
$\beta^{2}\langle$ $\big{(}\partial_{s_{o}^{1}}\tilde{H}-\partial_{s_{o}^{2}}\tilde{H}\big{)}\,\big{(}\partial_{s_{y}^{1}}\tilde{H}-\partial_{s_{y}^{2}}\tilde{H}\big{)}\rangle=-\beta\langle\big{(}\partial_{s_{y}^{1}}-\partial_{s_{y}^{2}}\big{)}\,\big{(}\partial_{s_{o}^{1}}\tilde{H}-\partial_{s_{o}^{2}}\tilde{H}\big{)}\rangle$
=$=\begin{cases}-\beta\,\langle\partial^{2}_{s_{o}^{1}}\tilde{H}+\partial^{2}_{s_{o}^{2}}\tilde{H}\big{)}\rangle&\mbox{ if $y=o$}\\ -\beta\langle\partial_{s_{y}^{1}}\partial_{s_{o}^{1}}\tilde{H}+\partial_{s_{y}^{2}}\partial_{s_{o}^{2}}\tilde{H}\big{)}\rangle&\mbox{ if $y\sim o$}\\ 0&\mbox{ otherwise}\\ \end{cases}$
=$=\begin{cases}\beta\,\sum\limits_{y\sim o}\Big{(}\langle e^{i(s_{o}^{1}-s_{y}^{1})}\rangle+\langle e^{-i(s_{o}^{2}-s_{y}^{2})}\rangle+2hr^{2}\langle\cos(rs_{o}^{1})\rangle\Big{)}&\mbox{ if $y=o$}\\ -\beta\big{(}\langle e^{i(s_{o}^{1}-s_{y}^{1})}\rangle+\langle e^{-i(s_{o}^{2}-s_{y}^{2})}\rangle\big{)}&\mbox{ if $y\sim o$}\\ 0&\mbox{ otherwise. }\\ \end{cases}$
This concludes the proof. ∎
6.1 Positivity of the Fourier transform
We next state an inequality which is a discrete analogue of the fact that the Fourier transform of a convex function is positive. The bound will be applied to two-point functions (4.12). Since our convexity property (4.13) fails at a neighbourhood of zero, we obtain a bound which may be negative, but which is negligible comparing to the volume $|\mathbb{S}_{L,K}|$. We define for any $M\in\mathbb{N}$, $\mathbb{Z}^{M}_{odd}:=(2\mathbb{Z}+1)\cap(-M/2,M/2]$ and $\mathbb{Z}^{M}_{even}:=(2\mathbb{Z})\cap(-M/2,M/2]$.
Lemma 6.2.
Let $G:\mathbb{S}_{L,K}\to\mathbb{R}$ be such that
  1. (i)
    $G(x)=G(-x)$ for any $x\in e_{1}\mathbb{Z}^{L}_{odd}\cup e_{2}\mathbb{Z}^{L}_{odd}\cup e_{3}\mathbb{Z}^{K}_{odd}$,
  2. (ii)
    $2G(x)-G(x-2e_{i})-G(x+2e_{i})\leq 0$ for any $x\in e_{i}\mathbb{Z}^{L}_{odd}\setminus\{\pm 1,\pm 3\}$ if $i\in\{1,2\}$ or for any $x\in e_{3}\mathbb{Z}^{K}_{odd}\setminus\{\pm 1,\pm 3\}$,
  3. (iii)
    $G(e_{i})\geq G(3e_{i})$ for each $i\in\{1,2,3\}$.
Then for every $k\in\mathbb{S}_{L,K}^{*}$ such that $k_{j}\not\in\{0,\pi\}$ for each $j\in\{1,2,3\}$ we have that
$\sum_{x\in e_{i}\mathbb{Z}_{odd}^{L}}G(x)\cos(k\cdot x)$ $\geq-(G(3e_{i})-G(5e_{i}))\frac{1-\cos 3k_{i}}{2\sin^{2}k_{i}},$ (6.3)
$\sum_{x\in e_{3}\mathbb{Z}_{odd}^{K}}G(x)\cos(k\cdot x)$ $\geq-(G(3e_{3})-G(5e_{3}))\frac{1-\cos 3k_{3}}{2\sin^{2}k_{3}},$ (6.4)
where $i\in\{1,2\}$.
Proof.
We note that for $i\in\{1,2\}$,
$\sum_{x\in e_{i}\mathbb{Z}^{L}_{odd}}G(x)\cos(k\cdot x)$ $=\frac{1}{(2\sin k_{i})^{2}}\sum_{x\in\mathbb{Z}^{L}_{odd}}(2G(x)-G(x-2e_{i})-G(x+2e_{i}))\cos(k\cdot x)$ (6.5)
$\sum_{x\in e_{3}\mathbb{Z}^{K}_{odd}}G(x)\cos(k\cdot x)$ $=\frac{1}{(2\sin k_{3})^{2}}\sum_{x\in\mathbb{Z}^{K}_{odd}}(2G(x)-G(x-2e_{3})-G(x+2e_{3}))\cos(k\cdot x).$ (6.6)
To see this, we notice that by the trigonometric identities
$\sum_{x\in e_{i}\mathbb{Z}^{L}_{odd}}G(x)\cos(k\cdot x)$ $=\sum_{x\in e_{i}\mathbb{Z}^{L}_{odd}}G(x)\frac{\sin\big{(}k\cdot(x+e_{i})\big{)}-\sin\big{(}k\cdot(x-e_{i})\big{)}}{2\sin(k_{i})}$
$=\frac{1}{2\sin k_{i}}\sum_{x\in e_{i}\mathbb{Z}^{L}_{even}}(G(x-e_{i})-G(x+e_{i}))\sin(k\cdot x)$
$=\frac{1}{2\sin k_{i}}\sum_{x\in e_{i}\mathbb{Z}^{L}_{even}}(G(x-e_{i})-G(x+e_{i}))\frac{\cos\big{(}k\cdot(x+e_{i})\big{)}-\cos\big{(}k\cdot(x-e_{i})\big{)}}{-2\sin(k_{i})}$
$=\frac{1}{(2\sin k_{i})^{2}}\sum_{x\in e_{i}\mathbb{Z}^{L}_{odd}}(2G(x)-G(x-2e_{i})-G(x+2e_{i}))\cos(k\cdot x),$
and similar computations lead to (6.6). By our assumption (ii) we can lower bound the Fourier sum by replacing $\cos(k\cdot x)$ by $1$ for any $x$ in the sum such that $|x|_{1}>3$. This leads to several cancellations in the sum and implies
$\sum_{x\in e_{i}\mathbb{Z}^{L}_{odd}}G(x)\cos(k\cdot x)$ $\geq\frac{1}{(2\sin k_{1})^{2}}\big{(}2(G(5e_{i})-G(3e_{i}))+2(2G(3e_{i})-G(e_{i})-G(5e_{i}))\cos(3k_{i})$
$+2(G(e_{i})-G(3e_{i}))\cos(k_{i})\big{)}$
$=\frac{1}{2(\sin k_{1})^{2}}\left((G(e_{i})-G(3e_{i}))(\cos k_{i}-\cos 3k_{i})+(G(3e_{i})-G(5e_{i}))(\cos 3k_{i}-1)\right),$
where we also used the condition (i). The previous expression and (iii) together lead to (6.3) as desired. Using analogous computations we obtain (6.4) and conclude the proof. ∎
6.2 Statement and proof of the Cauchy-Schwarz inequality
We now state a Cauchy-Schwarz inequality for the quantities $A$ and $B$ which have been defined above. Our inequality holds for small external magnetic fields, namely whose intensity vanishes proportionally to the volume. It will turn out that the error term, corresponding to the second term in the RHS of (6.7) below, is of smaller order if compared to the remaining terms.
Define the positive cone of the dual torus
$\mathbb{S}_{K,L,+}^{*}:=\big{\{}2\pi(\frac{n_{1}}{L},\frac{n_{2}}{L},\frac{n_{3}}{K})\in\mathbb{R}^{3}\,\,:\,\,n_{i}\in(-\frac{L}{4},\frac{L}{4}]\cap\mathbb{Z}\,,\text{ for }i=1,2\,,n_{3}\in(-\frac{K}{4},\frac{K}{4}]\cap\mathbb{Z}\big{\}}$
Theorem 6.3 (Cauchy-Schwarz inequality).
For any $\beta>0$ and $r\in\{1,2\}$ there exists $\tilde{h}_{0}>0$ and $c=c(\beta)\in\mathbb{R}$ such that, for any $\tilde{h}\in[0,\tilde{h}_{0})$, $L,K\in 2\mathbb{N}\cup\{1\}$ and $k\in\mathbb{S}_{K,L,+}^{*}$, we have that,
$\Big{(}{\langle A\circ B\rangle_{L,K,\beta,h_{L,K},r}}\Big{)}^{2}\leq\big{(}\langle A\circ A\rangle_{L,K,\beta,h_{L,K},r}+c\big{)}\langle B\circ B\rangle_{L,K,\beta,h_{L,K},r}$ (6.7)
where $h_{L,K}:=\frac{\tilde{h}}{|\mathbb{S}_{L,K}|}$.
Proof.
[…]
$\langle(\mu B-A)\circ(\mu B-A)\rangle$ $=\sum_{y\in{\mathcal{H}}}\cos(k\cdot y)\,\,\langle(\mu B_{o}-A_{o})(\mu B_{y}-A_{y})\rangle$
$+\sum_{y\in{\mathcal{V}}}\cos(k\cdot y)\,\,\langle(\mu B_{o}-A_{o})(\mu B_{y}-A_{y})\rangle$
$+\sum_{y\in\mathcal{W}}\cos(k\cdot y)\,\,\langle(\mu B_{o}-A_{o})(\mu B_{y}-A_{y})\rangle$
$+\langle(\mu B_{o}-A_{o})^{2}\rangle.$
For the first two terms on the right side above we first note that $(\mu B_{x}-A_{x})_{x\in\mathbb{S}_{L,K}}$ is a reflection invariant vector of functions. Hence by Propositions 4.3 and 4.4 the assumptions in Lemma 6.2 are satisfied for the two point function $G(y)=\langle(\mu B_{o}-A_{o})(\mu B_{y}-A_{y})\rangle$, and we can then apply that lemma to conclude that
$\sum_{y\in{\mathcal{H}}}\cos(k\cdot y)\langle(\mu B_{o}-A_{o})(\mu B_{y}-A_{y})\rangle+\sum_{y\in{\mathcal{V}}}\cos(k\cdot y)\langle(\mu B_{o}-A_{o})(\mu B_{y}-A_{y})\rangle$
$+\sum_{y\in\mathcal{W}}\cos(k\cdot y)\langle(\mu B_{o}-A_{o})(\mu B_{y}-A_{y})\rangle$
$\geq\sum_{i=1}^{3}(G(5e_{i})-G(3e_{i}))\cdot\frac{1-\cos 3k_{i}}{2\sin^{2}k_{i}}\geq-c,$
by our assumptions on $k$. Note that for the last inequality we observed that, for any $x\neq o$, we have that $G(x)=\langle\sin(rs_{o}^{1})\sin(rs_{x}^{1})\rangle$, and used the positive goodness condition and Lemma 5.1 to upper bound $G(x)\leq$ $\langle\cos(rs_{o}^{1})\cos(rs_{x}^{1})\rangle$ $\leq c$ for any $x\in\mathbb{S}_{L,K}$ uniformly in $L$.
For the diagonal term, using integration by parts for the second term in the expansion of the square, we obtain that
$\langle(\mu B_{o}-A_{o})^{2}\rangle=\langle\sin^{2}(rs^{1}_{o})\rangle-2\,\mu\beta r\,m+\mu^{2}\beta^{2}\langle(\partial_{s_{o}^{1}}\tilde{H}-\partial_{s_{o}^{2}}\tilde{H})^{2}\rangle.$
Using $\sin^{2}x=\frac{1}{2}-\frac{1}{2}\cos(2x)$ we obtain that the discriminant $\Delta$ satisfies,
$\frac{1}{4}\Delta$ $=\beta^{2}r^{2}m^{2}\,-\,\beta^{2}\frac{1}{2}\langle(\partial_{s_{o}^{1}}\tilde{H}-\partial_{s_{o}^{2}}\tilde{H})^{2}\rangle+\beta^{2}\frac{1}{2}\langle\cos(2rs^{1}_{o})\,\rangle\langle(\partial_{s_{o}^{1}}\tilde{H}-\partial_{s_{o}^{2}}\tilde{H})^{2}\rangle.$ (6.8)
Using again integration by parts and applying Proposition 5.4 we obtain that, for any $\beta>0$, if $\tilde{h}>0$ is not too large (depending on $\beta$), then there exists $c=c(\beta,\tilde{h})>0$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$,
$\beta^{2}\langle(\partial_{s_{o}^{1}}\tilde{H}-\partial_{s_{o}^{2}}\tilde{H})^{2}\rangle$ $=-\beta\,\langle\,\partial_{s_{o}^{1}}^{2}\tilde{H}\,\rangle-\beta\,\langle\,\partial_{s_{o}^{2}}^{2}\tilde{H}\,\rangle$
$=\beta\sum\limits_{y\sim o}\langle e^{i(s_{o}^{1}-s_{y}^{1})}\rangle+\beta\sum\limits_{y\sim o}\langle e^{-i(s_{o}^{2}-s_{y}^{2})}\rangle+2\beta hr^{2}\langle\cos(rs_{o}^{1})\rangle$
$\geq c,$
where for the last inequality we also used the lower bound in Lemma 5.3 and the positive goodness for $\cos(rs_{o}^{1})$. Using again Proposition 5.4 for the three terms in (6.8) we then deduce that, under the same assumptions, $\frac{1}{4}\Delta\leq-\frac{c}{2}<0$ where $c$ is the same constant as in the previous display. Combining the estimates for the diagonal and the non-diagonal terms we then obtain that, under the assumptions of our theorem,
$\forall\mu\in\mathbb{R}\quad\langle(\mu B-A)\circ(\mu B-A)\rangle\geq-c$
where $c=c(\beta)$ does not depend on $\mu$. We conclude the proposition by taking $\mu=\frac{\langle A\circ B\rangle}{\langle B\circ B\rangle}$ and using the fact that, by Lemma 6.1, $\langle A\circ B\rangle=\langle B\circ A\rangle$ and that $\langle B\circ B\rangle\geq 0$. ∎
6.3 Estimates for $\langle A\circ A\rangle$
In this section we provide some estimates for the sum over Fourier modes of $\langle A\circ A\rangle_{L,K,\beta,h,r}$.
Lemma 6.4.
There exists $c=c(\beta)<\infty$ such that, for any $h\in[0,1]$, $r\in\{1,2\}$, $K,L\in 2\mathbb{N}\cup\{1\}$, we have that
$\sum_{k\in\mathbb{S}_{L,K}^{*}}\langle A\circ A\rangle_{L,K,\beta,h,r}\leq c\,K\,L^{2}.$
Proof.
We omit the subscripts in the expectation. First note the trivial bound $\sum_{k\in\mathbb{S}_{L,K}^{*}}\langle\sin^{2}(rs_{o}^{1})\rangle\leq\,|\mathbb{S}_{L,K}^{*}|\,\langle\sin^{2}(rs_{o}^{1})\rangle\leq cKL^{2}.$ Here we used the trigonometric identity $\sin^{2}x=\frac{1}{2}-\frac{1}{2}\cos(2x)$, and $\langle\cos(2rs_{o}^{1})\rangle\geq 0$ by the positive goodness condition. We now upper bound the off-diagonal term. For any $x\in\mathbb{S}_{L,K}$ we define the sets ${\mathcal{H}}_{x}=\mathcal{H}+x$, ${\mathcal{V}}_{x}=\mathcal{V}+x$, ${\mathcal{W}}_{x}=\mathcal{W}+x$, where the sum is with respect to the torus metric, we also recall the definitions $\mathbb{Z}^{M}_{odd}=(2\mathbb{Z}+1)\cap(-M/2,M/2]$, $\mathbb{Z}^{M}_{even}=2\mathbb{Z}\cap(-M/2,M/2]$, and define $\mathbb{Z}^{M}:=(-M/2,M/2]\cap\mathbb{Z}$ for any $M\in\mathbb{N}$. Using translation invariance we have that
$KL^{2}\Biggl{(}\sum_{y\in{\mathcal{H}}}\cos(k\cdot y)\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle+\sum_{y\in{\mathcal{V}}}\cos(k\cdot y)\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle+\sum_{y\in\mathcal{W}}\cos(k\cdot y)\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle\Biggr{)}$
$=\sum_{x\in\mathbb{S}_{L,K}}\sum_{y\in{\mathcal{H}}_{x}}\cos(k\cdot(x-y))\langle\sin(rs_{x}^{1})\sin(rs_{y}^{1})\rangle+\sum_{x\in\mathbb{S}_{L,K}}\sum_{y\in{\mathcal{V}}_{x}}\cos(k\cdot(x-y))\langle\sin(rs_{x}^{1})\sin(rs_{y}^{1})\rangle$
$+\sum_{x\in\mathbb{S}_{L,K}}\sum_{y\in\mathcal{W}_{x}}\cos(k\cdot(x-y))\langle\sin(rs_{x}^{1})\sin(rs_{y}^{1})\rangle$
$\leq 2KL\sum_{x=(-\frac{L}{2}+1)e_{1}}^{\frac{L}{2}e_{1}}\sum_{y\in\mathbb{Z}^{L}_{odd}e_{1}+x}\cos(k\cdot(x-y))\langle\sin(rs_{x}^{1})\sin(rs_{y}^{1})\rangle$
$+L^{2}\sum_{x=(-\frac{K}{2}+1)e_{3}}^{\frac{K}{2}e_{3}}\sum_{y\in\mathbb{Z}^{K}_{odd}e_{3}+x}\cos(k\cdot(x-y))\langle\sin(rs_{x}^{1})\sin(rs_{y}^{1})\rangle$
$=2KL(\hat{G}^{1}_{\mathbb{Z}^{L}}(k)-\hat{G}^{1}_{\mathbb{Z}^{L}_{even}}(k)-\hat{G}^{1}_{\mathbb{Z}^{L}_{odd}}(k))+L^{2}(\hat{G}^{3}_{\mathbb{Z}^{K}}(k)-\hat{G}^{3}_{\mathbb{Z}^{K}_{even}}(k)-\hat{G}^{3}_{\mathbb{Z}^{K}_{odd}}(k)),$
where we defined for each $i\in\{1,2,3\}$, each $M\in\mathbb{N}$, and each set $A\in\{\mathbb{Z}^{M}_{odd}$, $\mathbb{Z}^{M}_{even},\mathbb{Z}^{M}\}$ the quantity
$\hat{G}^{i}_{A}(k)$ $:=\sum_{x,y\in e_{i}A}\cos(k\cdot(x-y))\langle\sin(rs_{x}^{1})\sin(rs_{y}^{1})\rangle.$
We now observe that
$\sum_{k\in\mathbb{S}_{L,K}^{*}}\hat{G}^{1}_{\mathbb{Z}^{L}}(k)=\sum_{k\in\mathbb{S}_{L,K}^{*}}\sum_{x,y=(-\frac{L}{2}+1)e_{1}}^{\frac{L}{2}e_{1}}\exp(ik\cdot(x-y))\langle\sin(rs_{x}^{1})\sin(rs_{y}^{1})\rangle\\ =|\mathbb{S}_{L,K}|\sum_{x=(-\frac{L}{2}+1)e_{1}}^{\frac{L}{2}e_{1}}\langle\sin^{2}(rs_{x}^{1})\rangle=KL^{3}\langle\sin^{2}(rs_{o}^{1})\rangle\leq CKL^{3},$
where for the second equality we used the Parseval identity and for the inequality used the fact that $\langle\sin^{2}(rs_{o}^{1})\rangle\leq\frac{1}{2}$, as observed before. Similarly, we deduce that $\sum_{k\in\mathbb{S}_{L,K}^{*}}\hat{G}^{1}_{\mathbb{Z}^{L}_{odd}}(k)\leq CKL^{3}$, $\sum_{k\in\mathbb{S}_{L,K}^{*}}\hat{G}^{1}_{\mathbb{Z}^{L}_{even}}(k)\leq CKL^{3}$, $\sum_{k\in\mathbb{S}_{L,K}^{*}}\hat{G}^{3}_{\mathbb{Z}^{K}}(k)\leq CK^{2}L^{2}$, $\sum_{k\in\mathbb{S}_{L,K}^{*}}\hat{G}^{3}_{\mathbb{Z}^{K}_{odd}}(k)\leq CK^{2}L^{2}$,
$\sum_{k\in\mathbb{S}_{L,K}^{*}}\hat{G}_{\mathbb{Z}^{K}_{even}}(k)\leq CK^{2}L^{2}$. Putting together all the previous bounds we obtain that
$\sum_{k\in\mathbb{S}_{L,K}^{*}}\biggl{(}\sum_{y\in{\mathcal{H}}}\cos(k\cdot y)\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle$ $+\sum_{y\in{\mathcal{V}}}\cos(k\cdot y)\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle$
$+\sum_{y\in\mathcal{W}}\cos(k\cdot y)\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle\biggr{)}$
$\quad\leq C^{\prime}KL^{2},$
and we conclude the lemma. ∎
We also need a lower bound for the large momentum region.
Lemma 6.5.
There exists $c=c(\beta)<\infty$ and $h_{0}>0$ such that, for any $\tilde{h}\in[0,h_{0}]$, $r\in\{1,2\}$, $K,L\in 2\mathbb{N}\cup\{1\}$,
$\sum_{k\in\mathbb{S}_{L,K}^{*}\setminus\mathbb{S}_{L,K,+}^{*}}\langle A\circ A\rangle_{L,K,\beta,h,_{L,K},r}\geq-c\,K\,L^{2}$
where $h_{L,K}=\frac{\tilde{h}}{|\mathbb{S}_{L,K}|}$.
Proof.
Recall that
$\langle A\circ A\rangle=\sum_{y\in\mathbb{Z}^{L}_{odd}e_{1}}\cos(k\cdot y)\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle+\sum_{y\in\mathbb{Z}^{L}_{odd}e_{2}}\cos(k\cdot y)\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle\\ +\sum_{y\in\mathbb{Z}^{K}_{odd}e_{3}}\cos(k\cdot y)\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle+\langle\sin^{2}(rs_{o}^{1})\rangle.$ (6.9)
To begin, we sum over $k$ and obtain for any $y\in\mathbb{Z}$, $i\in\{1,2\}$,
$\sum_{k\in\mathbb{S}_{L,K}^{*}\setminus\mathbb{S}_{L,K,+}^{*}}\cos(k_{i}\,y)=\sum_{n_{3}\in((-\frac{K}{2},\frac{K}{2}]\setminus(-\frac{K}{4},\frac{K}{4}])\cap\mathbb{Z}}\sum_{n_{1},n_{2}\in((-\frac{L}{2},\frac{L}{2}]\setminus(-\frac{L}{4},\frac{L}{4}])\cap\mathbb{Z}}\cos(\frac{2\pi n_{1}}{L}y)$
$=\frac{1}{4}KL\sum_{n_{1}\in((-\frac{L}{2},\frac{L}{2}]\setminus(-\frac{L}{4},\frac{L}{4}])\cap\mathbb{Z}}\cos(\frac{2\pi n_{1}}{L}y)$ (6.10)
Note that
$\sum_{n_{1}\in([-\frac{L}{2},\frac{L}{2}]\setminus[-\frac{L}{4},\frac{L}{4}])\cap\mathbb{Z}}\cos(\frac{2\pi n_{1}}{L}y)$ $=\sum_{n_{1}\in([-\frac{L}{2},\frac{L}{2}]\setminus[-\frac{L}{4},\frac{L}{4}])\cap\mathbb{Z}}e^{i\frac{2\pi n_{1}}{L}y}$
$=e^{i\frac{\pi}{2}y}e^{i\frac{2\pi}{L}y}\frac{1-e^{i\frac{\pi}{2}y}}{1-e^{i\frac{2\pi}{L}y}}+e^{-i\frac{\pi}{2}y}e^{-i\frac{2\pi}{L}y}\frac{1-e^{-i\frac{\pi}{2}y}}{1-e^{-i\frac{2\pi}{L}y}}$
$=\frac{i^{y}(1-i^{y})(e^{i\frac{2\pi}{L}y}-1)+(-i)^{y}(1-(-i)^{y})(e^{-i\frac{2\pi}{L}y}-1)}{2-2\cos(\frac{2\pi}{L}y)}$
Since $y$ is odd, we use $i^{2y}=-1$ to simplify the summation above as
$\sum_{n_{1}\in([-\frac{L}{2},\frac{L}{2}]\setminus[-\frac{L}{4},\frac{L}{4}])\cap\mathbb{Z}}\cos(\frac{2\pi n_{1}}{L}y)=\frac{i^{y+1}\sin(\frac{2\pi}{L}y)}{1-\cos(\frac{2\pi}{L}y)}-1.$
By comparing the previous expression with (6.3) we see that there are mismatch terms at $n_{1}=-\frac{L}{2}$ and $n_{1}=-\frac{L}{4}$, concluding that
$\sum_{k\in\mathbb{S}_{L,K}^{*}\setminus\mathbb{S}_{L,K,+}^{*}}\cos(k_{i}y)=\frac{1}{4}KL\frac{i^{y+1}\sin(\frac{2\pi}{L}y)}{1-\cos(\frac{2\pi}{L}y)}+O(KL),$ (6.11)
where the last term does not depend on $y$. We can now lower bound the sum over $k$ of $\langle A\circ A\rangle$ to prove the lemma. By (5.9), we see that $\langle\sin^{2}(rs_{o}^{1})\rangle=\frac{1}{2}-\frac{1}{2}\langle\cos(2rs_{o}^{1})\rangle\geq\frac{1}{2}-ch_{0}$. This implies, $\langle\sin^{2}(rs_{o}^{1})\rangle>0$ by taking $h_{0}$ sufficiently small. Using this observation, (6.9) and (6.11) we obtain that
$\sum_{k\in\mathbb{S}_{L,K}^{*}\setminus\mathbb{S}_{L,K,+}^{*}}\langle A\circ A\rangle$ $\geq\sum_{y\in\mathbb{Z}^{L}_{odd}e_{1}}\frac{1}{4}KL\frac{i^{y+1}\sin(\frac{2\pi}{L}y)}{1-\cos(\frac{2\pi}{L}y)}\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle$
$+\sum_{y\in\mathbb{Z}^{L}_{odd}e_{2}}\frac{1}{4}KL\frac{i^{y+1}\sin(\frac{2\pi}{L}y)}{1-\cos(\frac{2\pi}{L}y)}\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle$
$+\sum_{k\in\mathbb{S}_{L,K}^{*}\setminus\mathbb{S}_{L,K,+}^{*}}\sum_{y\in\mathbb{Z}^{K}_{odd}e_{3}}\cos(k\cdot y)\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle$
$-cKL\Biggl{(}\sum_{y\in\mathbb{Z}^{L}_{odd}e_{1}}|\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle|+\sum_{y\in\mathbb{Z}^{L}_{odd}e_{2}}|\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle|\Biggr{)}.$ (6.12)
Since $\{\frac{i^{y+1}\sin(\frac{2\pi}{L}y)}{1-\cos(\frac{2\pi}{L}y)}\}$ is an alternating sequence with decreasing magnitude in $(0,\pi)\cap(2\mathbb{N}+1)$, together with the monotonicity of the two point function, $\langle\sin(rs_{o}^{1})\sin(rs_{(2n-1)e_{i}}^{1})\rangle\geq$ $\langle\sin(rs_{o}^{1})\sin(rs_{(2n+1)e_{i}}^{1})\rangle$ for odd integers $2n-1\in(1,L/2)$, and symmetry we see that the first two sums on the right side of (6.3) are bounded from below by the sum of the $y=e_{1}$ and $y=-e_{1}$ term:
$\ \sum_{y\in\mathbb{Z}^{L}_{odd}e_{1}}\frac{1}{4}KL\frac{i^{y+1}\sin(\frac{2\pi}{L}y)}{1-\cos(\frac{2\pi}{L}y)}\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle+\sum_{y\in\mathbb{Z}^{L}_{odd}e_{2}}\frac{1}{4}KL\frac{i^{y+1}\sin(\frac{2\pi}{L}y)}{1-\cos(\frac{2\pi}{L}y)}\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle$
$\geq-cKL\frac{L}{2\pi}(\langle\sin(rs_{o}^{1})\sin(rs_{e_{1}}^{1})\rangle+\langle\sin(rs_{o}^{1})\sin(rs_{e_{2}}^{1})\rangle)\geq-cKL^{2}.$ (6.13)
On the other hand, using the monotonicity of the two point function we have
$-cKL\Biggl{(}\sum_{y\in\mathbb{Z}^{L}_{odd}e_{1}}\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle+\sum_{y\in\mathbb{Z}_{odd}e_{2}}\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle\Biggr{)}\\ \geq-cKL^{2}(\langle\sin(rs_{o}^{1})\sin(rs_{e_{1}}^{1})\rangle+\langle\sin(rs_{o}^{1})\sin(rs_{e_{2}}^{1})\rangle)\geq-c^{\prime}KL^{2},$ (6.14)
where for the last inequality we use Lemma 5.1 to deduce that $\langle\sin(rs_{o}^{1})\sin(rs_{e_{1}}^{1})\rangle\leq c^{2}$ uniformly in $L$ and $K$.
For the sum of $y$ in $e_{3}$ direction, the same computation as (6.3) and (6.11) yields
$\sum_{k\in\mathbb{S}_{L,K}^{*}\setminus\mathbb{S}_{L,K,+}^{*}}\cos(k_{3}\,y)=\frac{1}{4}L^{2}\sum_{n_{3}\in((-\frac{K}{2},\frac{K}{2}]\setminus(-\frac{K}{4},\frac{K}{4}])\cap\mathbb{Z}}\cos(\frac{2\pi n_{3}}{K}y)=\frac{1}{4}L^{2}\frac{i^{y+1}\sin(\frac{2\pi}{K}y)}{1-\cos(\frac{2\pi}{K}y)}+O(L^{2}),$
where the last term in the RHS does not depend on $K$ or $y$. Using that $\{\frac{i^{y+1}\sin(\frac{2\pi}{K}y)}{1-\cos(\frac{2\pi}{K}y)}\}$ is an sequence series with decreasing magnitude in $(0,K/2)\cap(2\mathbb{N}+1)$, the monotonicity of the two point function, symmetry and Lemma 5.1, as before, we obtain that
$\sum_{k\in\mathbb{S}_{L,K}^{*}\setminus\mathbb{S}_{L,K,+}^{*}}\sum_{y\in\mathbb{Z}^{K}_{odd}e_{3}}|\cos(k\cdot y)\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle|\\ \geq-cL^{2}\frac{K}{2\pi}\langle\sin(rs_{o}^{1})\sin(rs_{e_{3}}^{1})\rangle-cKL^{2}\langle\sin(rs_{o}^{1})\sin(rs_{e_{3}}^{1})\rangle\geq-c^{\prime}KL^{2}.$
Combing with (6.13) and (6.14) above we conclude the proof of the lemma. ∎
7 Mermin-Wagner Theorem
We deduce our Mermin Wagner theorem from the following upper bound for the magnetisation.
Proposition 7.1 (Magnetisation bound).
Suppose that $\beta\geq 0$, $r\in\{1,2\}$, and that $\tilde{h}>0$ is small enough. There a constant $c=c(\beta)<\infty$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$ satisfying $K\leq(\log L)^{\frac{1}{2}}$, we have
$m_{\mathbb{S}_{L,K}}(\beta,h_{L,K},r)\leq c\sqrt{\frac{K}{\log L}},$
where $h_{L,K}:=\frac{\tilde{h}}{|\mathbb{S}_{L,K}|}$.
Proof.
From (6.7) and Lemma 6.1 we deduce that
$\frac{r^{2}\,m_{\mathbb{S}_{L,K}}(\beta,h_{L,K},r)^{2}}{\frac{2}{\beta}|k|^{2}+8h_{L,K}\,m_{\mathbb{S}_{L,K}}(\beta,h_{L,K},r)}\\ \leq\frac{r^{2}\,m_{\mathbb{S}_{L,K}}(\beta,h_{L,K},r)^{2}}{\sum\limits_{z\sim o}p_{z}\,\big{(}1-\cos(k\cdot z)\big{)}+2r^{2}\,h_{L,K}\,m_{\mathbb{S}_{L,K}}(\beta,h_{L,K},r)}\\ \leq\big{(}\langle A\circ A\rangle_{L,K,\beta,h_{L,K},r}+c\big{)}.$
In the sequel we omit the dependences of $m_{\mathbb{S}_{L,K}}$, which continues to depend on $\beta$ and $h_{L,K}$, and $r$, in order to simplify the notation.
Summing over $k\in\mathbb{S}_{L,K,+}^{*}$, applying Lemma 6.4 and 6.5 to upper bound the RHS of the previous expression, and dividing by $|\mathbb{S}_{L,K}|$ leads to
$\frac{1}{|\mathbb{S}_{L,K}|}\sum_{k\in\mathbb{S}_{L,K,+}^{*}}\frac{m^{2}}{\frac{2}{\beta}|k|^{2}+\frac{8\tilde{h}}{|\mathbb{S}_{L,K}|}m}\leq c,$ (7.1)
which holds for any $L,K\in 2\mathbb{N}\cup 1$.
Let us now define the set $\mathcal{U}:=\{(L,K)\in(2\mathbb{N}+1)^{2}\,:\,K\leq\sqrt{\log L}\}$. Suppose that for any $\varphi>0$ we have that
$m_{\mathbb{S}_{L,K}}(\beta,h_{L,K},r)\geq\varphi\sqrt{\frac{K}{\log L}}\,\,\mbox{ for infinitely many choices of $(L,K)\in\mathcal{U}$.}$ (7.2)
We will find a contradiction with this assumption, deducing that there exists a value $\varphi>0$ such that $m\geq\varphi\sqrt{\frac{K}{\log L}}$ can only hold for finitely many choices of $(L,K)\in\mathcal{U}$. This implies the claim of the proposition, as desired.
Let us now show that (7.2) leads to a contradiction. From (7.1) and (7.2) we deduce that, for infinitely many choices of $(L,K)\in\mathcal{U}$,
$\frac{1}{|\mathbb{S}_{L,K}|}\sum_{k\in\mathbb{S}_{L,K,+}^{*}}\frac{\varphi^{2}}{c_{2}|k|^{2}+\frac{c_{1}}{|\mathbb{S}_{L,K}|}}\leq c_{3}\frac{\log L}{K},$ (7.3)
where we also applied Proposition 5.4 in order to upper bound $m$ in the denominator. We now lower bound the term in the LHS of (7.3) from the term below,
$\frac{1}{|\mathbb{S}_{L,K}|}\sum_{\begin{subarray}{c}k\in\mathbb{S}_{L,K,+}^{*}:\\ |k|^{2}\geq\frac{c_{1}}{c_{2}}\frac{1}{|\mathbb{S}_{L,K}|}\end{subarray}}\frac{\varphi^{2}}{2c_{2}|k|^{2}}+\sum_{\begin{subarray}{c}k\in\mathbb{S}_{L,K,+}^{*}:\\ |k|^{2}<\frac{c_{1}}{c_{2}}\frac{1}{|\mathbb{S}_{L,K}|}\end{subarray}}\frac{\varphi^{2}}{2c_{1}}.$
The main contribution comes from the first sum, which, after reducing the sum to all $k$ such that $k_{3}=0$ and using the change of variable $j_{i}=\frac{Lk_{i}}{2\pi}$ for $i=1,2$, equals
$\frac{1}{8\pi^{2}c_{2}}\frac{1}{K}\sum_{\begin{subarray}{c}j_{1},j_{2}\in\{-\frac{L}{4}+1,-\frac{L}{4}+2,\ldots\frac{L}{4}\}:\\ j_{1}^{2}+j_{2}^{2}\geq c/K\end{subarray}}\frac{\varphi^{2}}{(j_{1}^{2}+j_{2}^{2})}.$
We finally note that the previous term is bounded from below by $c\,\varphi^{2}\,\frac{1}{K}\,\log(L)$. Summarizing, we deduce that for any $\varphi>0$, and infinitely many values $(L,K)\in\mathcal{U}$
$c_{4}\varphi^{2}\frac{\log(L)}{K}\leq c_{3}\frac{\log L}{K},$ (7.4)
which is not possible. Hence, we obtained the desired contradiction and concluded the proof. ∎
We now state our main theorem in the probabilistic framework of random loops with arbitrary weight function.
Theorem 7.2 (Mermin-Wagner theorem for double dimers).
Let $\beta\geq 0$, $\ell\in\{1,2\}$ be arbitrary. There exists $c=c(\beta)<\infty$ such that for any $L,K\in 2\mathbb{N}\cup\{1\}$ satisfying $K\leq\sqrt{\log(L)}$ we have
$\sum_{x\in\mathbb{S}_{L,K}^{o}}\frac{\mathcal{G}_{L,K,\beta}^{(1)}(o,x)}{|\mathbb{S}^{o}_{L,K}|}$ $\leq c\sqrt{\frac{K}{\log L}}$ (7.5)
$\sum_{x\in\mathbb{S}_{L,K}}\frac{\mathcal{G}_{L,K,\beta}^{(\ell)}(o,x)}{|\mathbb{S}_{L,K}|}$ $\leq c\sqrt{\frac{K}{\log L}},$ (7.6)
where the second inequality holds only if $U$ is nice when $\ell=1$ and only if $U$ is fast decaying when $\ell=2$.
We may conclude the proof of the Mermin Wagner theorem.
Proof of Theorem 1.1 when $N=1$ and of Theorems 1.2, 2.1, and 7.2.
Fix $\beta>0$, $\tilde{h}>0$ small enough. We apply the first claim in Proposition 5.4 and Proposition 7.1 to deduce our main theorem for the spin system, Theorem 2.1.
We now apply Proposition 3.3 and the observation $\sum_{x\in\mathbb{S}_{L,K}}\mathcal{G}_{L,K,\beta}^{(2)}(o,x)=2\sum_{x\in\mathbb{S}_{L,K}^{o}}\mathcal{G}_{L,K,\beta}^{(2)}(o,x)$ to deduce from Theorem 2.1 the proof of Theorem 7.2.
Since for the monomer double dimer model $U$ is fast decaying (and nice), when $\beta>0$ is finite we can apply Theorem 7.2 with $\ell=2$ and (3.6) to deduce our main theorem on the expected length of the loops in the monomer double dimer model, Theorem 1.1, when $N=1$ and $\rho=1/\beta>0$.
Moreover, we use (7.5) (which does not require the assumption that $U$ is nice) for $U=\mathbbm{1}_{\{n=1\}}$, corresponding to the double dimer model, (3.7), and the fact that $\mathcal{C}(o,x)=0$ if $x\in\mathbb{S}_{L,K}^{e}$ to obtain our second main theorem on monomer correlations in the dimer model, Theorem 1.2.
In order to obtain Theorem 1.1 when $N=1$ and $\rho=0$ we use that
$\mathbb{P}_{L,K,1,0}(x\leftrightarrow y)=\sum\limits_{z\sim x,k\sim y}\,\mathcal{C}(x,y)\mathcal{C}(z,k)\leq d^{2}\,\mathcal{C}(x,y),$
where the identity follows from the fact that by [21, Figure 2] $\mathcal{C}(x,y)\mathcal{C}(z,k)$ corresponds to the probability that in the double dimer model a loop connects $x$ to $y$ with the blue dimers of such loop touching $x$ and $y$ being on the edges $\{x,z\}$ and $\{y,k\}$, while the inequality follows from the fact that $\mathcal{C}(z,k)\leq 1$ for any $z,k\in\mathbb{S}_{L,K}$ by [23, equation (2.25)]. Hence using the previous inequality and Theorem 1.2 we deduce Theorem 1.1 when $\rho=0$ and $N=1$.
7.1 Multiple dimer model
In order to treat the case of integer $N>1$ we need to generalise the spin representation considering $S=[0,2\pi)^{2N}$ as a local spin space. For each $\boldsymbol{s}\in S^{V}$, the Hamiltonian can be generalised to,
$H(\boldsymbol{s})=\sum\limits_{u=1}^{2N}\sum\limits_{\{i,j\}\in E}e^{i(s_{i}^{u}-s_{j}^{u})}+\,h\sum\limits_{k\in V}\,2\,\cos(rs^{1}_{k}),$
and the function $g$ has to be generalised as follows,
$g(s)=\frac{1}{(2\pi)^{2N}}\sum\limits_{n=(n_{1},\ldots n_{N})\in\mathbb{N}_{0}^{N}}U(n_{1},\dots,n_{N})e^{-\sum_{u=1}^{N}n_{u}\,i\,(s^{2u-1}-s^{2u})}$
where $U:\mathbb{N}^{N}\mapsto\mathbb{R}$ is the weight function and $s=(s^{1},\ldots,s^{N})$.
The double dimer representation which is obtained by applying the strategy of Proposition 8 below leads to a double dimer representation which is analogous to the one which has been introduced in Section 3, the only difference is that dimers have $2N$ possible colours, $1,\ldots,2N$, and at each vertex the number of dimers of odd colour $i\in[1,2N]$ equals the number of dimers of colour $i+1$. In this representation, each vertex touched by $n_{i}$ dimers of colour $2i$ (which is then also touched by $n_{i}$ dimers of colour $12i-$ 1) for each $i\in\{1,\ldots,N\}$ receives a multiplicative weight $U(n_{1},\ldots n_{N})$.
Pairings can be introduced analogously to Section 3 to obtain open and closed path, namely at each vertex $x\in V$ dimers of odd colour $i$ can only be paired to paths of colour $i+1$ and viceversa. This leads to a random walk loop soup in which each close path is a sequence of dimers whose colours are $2i$, $12i-$ 1, $2i$, $12i-$ 1, $\ldots$,… , or $12i-$ 1, $2i$, $12i-$ 1, $2i$, $\ldots$ for some integer $i\in[0,N]$.
We obtain the monomer double dimer-model (1.1) with arbitrary $N\in\mathbb{N}$ in the special case,
$U(n_{1},n_{2},\ldots n_{N})=\begin{cases}1&\mbox{ if $n_{1}+\ldots+n_{N}\in\{0,1\}$,}\\ 0&\mbox{ otherwise.}\end{cases}$
Indeed, in this case the local time of each vertex is either zero or one, the same as in (1.1), and each loop of length two has multiplicity $N$ (we have $N$ possibilities for choosing whether the dimers constituting this loop have colours $1$ and $2$, $3$ and $4$, $\ldots$, or $12N-$ 1 and $N$) while each loop of length greater than two has multiplicity $2N$ (with each case corresponding to one of the possible sequences $2i$, $12i-$ 1, …$2i$ or $12i-$ 1, $2i$, …$12i-$ 1). This is in correspondence to (1.1), in which each loop of length two has multiplicity $1$ and each loop of length greater than two has multiplicity $2$ and, in addition, there is a multiplicative weight $N$ for each loop.
8 Proof of Proposition 3.1
Proof.
We observe that the factor $e^{\beta H(\boldsymbol{s})}$ in (2.4) can be seen as a product of exponential functions. Each such exponential function is associated to an edge of the extended graph and to a colour. Recall the definition of $G=(V,E)$ , of $G_{en}=(V_{en},E_{en})$, $E_{s}$, and $E_{g}$. Assume that $A\subset V$. We start by performing a Taylor expansion for each such exponential in the product, obtaining
$\int_{S^{V}}\boldsymbol{ds}\,\boldsymbol{g}(\boldsymbol{s})\prod_{z\in A}g_{u,z}(s_{z})\,\,e^{\beta H(\boldsymbol{s})}\,\prod_{x\in V^{e}}e^{i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2})}\prod_{x\in V^{o}}e^{-i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2})}\\ =\sum\limits_{m^{1}\in\mathbb{N}_{0}^{E}}\prod_{e\in E}\frac{\beta^{m^{1}_{e}}}{m^{1}_{e}!}\sum\limits_{m^{2}\in\mathbb{N}_{0}^{E}}\prod_{e\in E}\frac{\beta^{m^{2}_{e}}}{m^{2}_{e}!}\sum\limits_{m^{g}\in\mathbb{N}_{0}^{E_{g}}}\prod_{e\in E_{g}}\frac{(\beta h)^{m^{g}_{e}}}{m^{g}_{e}!}\\ \times\int_{S^{V}}\boldsymbol{ds}\,\boldsymbol{g}(\boldsymbol{s})\prod_{z\in A}g_{u,z}(s_{z})\,\,\prod_{x\in V^{e}}e^{i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2}+\,\tilde{n}_{x}(m^{1})s_{x}^{1}\,-\tilde{n}_{x}(m^{2})s_{x}^{2})}\prod_{x\in V^{o}}e^{-i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2}+\,\tilde{n}_{x}(m^{1})s_{x}^{1}\,-\tilde{n}_{x}(m^{2})s_{x}^{2})}\\ \times\prod_{x\in V}(2\cos(rs_{x}^{1}))^{m_{\{x,g\}}^{g}},$ (8.1)
where $\tilde{n}_{x}(m^{i})=\sum_{y\sim x}m_{\{x,y\}}^{i}$, recalling that $\sim$’ is the neighbourhood relation in $G$ and not in $G_{en}$. We interpret $m_{\{x,y\}}^{i}$ as the number of dimers of colour $i$ on $\{x,y\}$ and $m_{\{x,g\}}^{g}$ as the number of dimers on $\{x,g\}$ (independently from their colour). Given realisation of the vector $m^{g}$ in the previous sum, we perform the following expansions at each vertex
[…]
obtaining that the RHS of (8.1) equals
$\sum\limits_{m^{1}\in\mathbb{N}_{0}^{E}}\sum\limits_{m^{2}\in\mathbb{N}_{0}^{E}}\sum\limits_{m^{g,1}\in(r\mathbb{N}_{0})^{E_{g}}}\sum\limits_{m^{g,2}\in(r\mathbb{N}_{0})^{E_{g}}}\prod_{e\in E}\frac{\beta^{m^{1}_{e}}}{m^{1}_{e}!}\prod_{e\in E}\frac{\beta^{m^{2}_{e}}}{m^{2}_{e}!}\prod_{e\in E_{g}}\frac{(\beta h)^{\frac{m^{g,1}_{e}}{r}}}{\frac{m^{g,1}_{e}}{r}!}\prod_{e\in E_{g}}\frac{(\beta h)^{\frac{m^{g,2}_{e}}{r}}}{\frac{m^{g,2}_{e}}{r}!}\\ \times\int_{S^{V}}\boldsymbol{ds}\,\boldsymbol{g({s})}\prod_{z\in A}g_{u,z}(s_{z})\prod_{x\in V^{e}}e^{i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2}+\,\tilde{n}_{x}(m^{1})s_{x}^{1}\,-\tilde{n}_{x}(m^{2})s_{x}^{2}+m_{\{x,g\}}^{g,1}s_{x}^{1}-m_{\{x,g\}}^{g,2}s_{x}^{1})}\\ \prod_{x\in V^{o}}e^{-i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2}+\,\tilde{n}_{x}(m^{1})s_{x}^{1}\,-\tilde{n}_{x}(m^{2})s_{x}^{2}+m_{\{x,g\}}^{g,1}s_{x}^{1}-m_{\{x,g\}}^{g,2}s_{x}^{1})},$ (8.2)
where we set $m^{g,2}_{\{x,g\}}=m^{g}_{\{x,g\}}-m^{g,1}_{\{x,g\}}$, and interpret $m^{g,i}_{\{x,g\}}$ as the number of dimers of colour $i$ on $\{x,g\}$. Given $n_{1},n_{2}\in\mathbb{Z}$ we use for each vertex in $V\setminus A$ the identity
$(\frac{1}{2\pi})^{2}\int_{S}e^{i(n_{1}s^{1}-n_{2}s^{2})}\sum\limits_{n=0}^{\infty}U(n)e^{-i\,n\,(s^{1}-s^{2})}ds=U(n_{1})\mathbbm{1}_{\{n_{1}=n_{2}\}},$
and for each vertex in $A$ the identity
[…]
obtaining that, for any fixed $m^{1},m^{2},m^{1,g}$, $m^{2,g}$ in the sum (8.2), the corresponding integral equals
$\prod_{x\in V\setminus A}\mathbbm{1}_{\{\tilde{n}_{x}(m^{1})+m_{\{x,g\}}^{g,1}+u_{x}^{1}=\tilde{n}_{x}(m^{2})+u_{x}^{2}+m_{\{x,g\}}^{g,2}\}}U(\tilde{n}_{x}(m^{1})+u_{x}^{1}+m_{\{x,g\}}^{g,1}-m_{\{x,g\}}^{g,2})\\ \times\prod_{x\in A}\mathbbm{1}_{\{\tilde{n}_{x}(m^{1})+m_{\{x,g\}}^{g,1}+u_{x}^{1}-m_{\{x,g\}}^{g,2}=\tilde{n}_{x}(m^{2})+u_{x}^{2}=u\}}U(u).$ (8.3)
For $x\in V$, we further write $u_{x}^{i}=(u_{+}^{i})_{x}+(u_{-}^{i})_{x}$ (recalling the definitions in the statement of the proposition) and interpret $(u_{+}^{1})_{x}-(u_{-}^{2})_{x}$ as the number of $1$-dimers on $\{x,s\}$ and $-(u_{-}^{1})_{x}+(u_{+}^{2})_{x}$ as the number of $2$-dimers on $\{x,s\}$ (so that these numbers are always non negative). Thus, $n_{x}^{1}=n_{x}^{1}(m^{1},m^{1,g},u^{1},u^{2})=\tilde{n}_{x}(m^{1})+m_{\{x,g\}}^{1,g}+(u^{1}_{+})_{x}-(u^{2}_{-})_{x}$ corresponds to the number of $1$-dimers touching $x\in V$ and $n_{x}^{2}=n_{x}^{2}(m^{2},m^{2,g},u^{1},u^{2})=\tilde{n}^{2}_{x}(m^{2})+m_{x,g}^{2,g}-(u^{1}_{-})_{x}+(u^{2}_{+})_{x}$ corresponds to the number of $2$-dimers touching $x\in V$. We then obtain that, for any pair $m=(m^{1},m^{2})\in\mathbb{N}_{0}^{E_{en}}\times\mathbb{N}_{0}^{E_{en}}$, the previous expression can be reformulated as follows
$\prod_{x\in V\setminus A}\mathbbm{1}_{\{n_{x}^{1}(m)=n_{x}^{2}(m)\}}U\big{(}n^{1}_{x}(m)-\,m_{\{x,g\}}^{g,2}+(u_{-}^{1})_{x}+(u_{-}^{2})_{x}\big{)}\prod_{x\in A}\mathbbm{1}_{\{n_{x}^{1}(m)=n_{x}^{2}(m)=u+m_{\{x,g\}}^{g,2}-(u_{-}^{1})_{x}-(u_{-}^{2})_{x}\}}U(u).$ (8.4)
Recalling the definition of $\Sigma$ and the notation (3.1) we then obtain that the LHS of (8.1) equals
$\sum\limits_{\begin{subarray}{c}m=(m^{1},m^{2})\in\Sigma:\\ \partial m=(u^{1}_{+}-u^{2}_{-},\,u^{2}_{+}-u^{1}_{-})\\ n_{z}(m)=m_{\{z,g\}}^{2}-(u_{-}^{1})_{z}-(u_{-}^{2})_{z}+u\,\,\,\forall z\in A\end{subarray}}\tilde{\mu}_{\beta,h,r}(m)\,\,\prod_{x\in V}U\big{(}n_{x}(m)-\,m_{\{x,g\}}^{2}+(u_{-}^{1})_{x}+(u_{-}^{2})_{x}\big{)}.$
By further observing that for each $m\in\Sigma$ we have precisely $n_{x}(m)!$ possible pairings at each vertex $x\in V$, we deduce (3.9) and (3.10). We also see that if we replace $g_{u,z}$ with $g_{z}$ in the product in (8.1) we obtain that the sum over $m^{g}\in\mathbb{N}_{0}^{E_{g}}$ in the next expression replaced by a sum over $m^{g}\in\mathbb{N}_{0}^{E_{g}}$ such that $m_{\{x,g\}}=0$ for each $x\in A$ (i.e, no dimer on $\{x,g\}$ for $x\in A$ is present), by further assuming that $u_{x}^{i}=0$ for each $x\in V$ we then obtain that (8.4) is replaced by
$\prod_{x\in V\setminus A}\mathbbm{1}_{\{n_{x}^{1}(m)=n_{x}^{2}(m)\}}U\big{(}n^{1}_{x}(m)\big{)}\prod_{x\in A}\mathbbm{1}_{\{n_{x}^{1}(m)=n_{x}^{2}(m)=0\}}U(0),$ (8.5)
this leads to (3.11) following the same steps as for the previous display. ∎
Acknowledgements
L.T. thanks Volker Betz and Benjamin Lees for sharing discussions on a closely related complex spin representation. The authors thank Tom Spencer for illuminating comments and Alexandra Quitmann for producing the figure. L.T. thank the German Research Foundation (project number 444084038, priority program SPP2265) for financial support. W.W was partially supported by MOST grant 2021YFA1002700, NSFC grant 20220903 NYTP and SMEC grant 2010000080.